### v1

# Enhanced Single-Photon Detector: Achieving Superconducting-Level Performance with Conventional Quantum Technology

Abstract
High-performance single-photon detectors (SPDs) are indispensable components for quantum optical tasks. However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and challenging operational conditions (e.g., cryogenics), which hinder scalable commercial deployment. To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of low-technology SPDs. Utilizing only standard quantum optical components, the ESPD scheme transforms a legacy non-superconducting SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) $10^{-2}$, into a device with superior performance metrics, achieving DE higher than $95\%$ and DCR below $10^{-9}$. This level of performance is comparable to or surpasses recently designed superconducting SPDs, allowing the minimal tolerable channel transmission rate for Quantum Key Distribution (QKD) protocols to be reduced by several orders of magnitude. Furthermore, the scheme’s device requirements are moderate, relying on readily available current technology, which ensures near-term experimental feasibility. The ESPD framework thus provides a clear, scalable path toward the large-scale deployment of high-performance SPDs and the commercialization of quantum communication technologies.
keywords:
Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
1 Introduction
Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurement is indispensable for information readout across virtually all applications. The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former influences the efficiency of nearly all quantum operations, while the latter constitutes the ultimate limitation in high-attenuation scenarios. Regarding DE, numerous quantum experiments rely on SPD signals to retrieve results. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3]. In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where entangled pairs are generated and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16]. Regarding DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD). The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, which is significantly lower than 50% [17, 18, 19, 20]. However, communication channels induce exponential photon attenuation [21, 22, 23, 24]. If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics. Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
Given their critical importance, the development of SPDs has been a continuous research focus since the inception of quantum technologies. While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR. For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP detector presented in 2017 operates at 223 K with a DE of 27.5% and a DCR of $10^{-6}$ [29]. These specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$. To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These detectors achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude. This capability enables the implementation of demanding quantum tasks, such as long-distance QKD distance [19, 34, 35, 36]. A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
Year Researchers (et al.) Operation temperature Detection efficiency Dark count rate
2020 P. Hu   [30] 2.10 K 95.0 % $0.5\times 10^{-5}$
2021 J. Chang [31] 2.50 K 99.5 % $1.1\times 10^{-3}$
2021 G. Z. Xu [32] 0.84 K 92.2 % $3.6\times 10^{-5}$
2023 I. Craiciu [33] 0.90 K 78.0 % $1.0\times 10^{-7}$
2024 I. Charaev [37] 20.0 K 7.60 % $1.0\times 10^{-3}$
However, as illustrated in the table, the reliance on superconductors necessitates stringent operating conditions, specifically with extremely low temperatures. Furthermore, these systems often entail high technical complexity and cost, hindering scalable commercial deployment.
To address these limitations, we propose a design for SPDs that meets advanced performance requirements, and even potentially outperforms superconducting devices, while relying on accessible, non-superconducting technology suitable for room-temperature operation. The proposed method, termed the Enhanced SPD (ESPD) scheme, iteratively upgrades the performance of an existing SPD level by level. As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] to one achieving a DE exceeding 95% and a DCR below $10^{-9}$, thereby rivaling state-of-the-art devices. Our method not only provides a high-performance SPD design realizable with current (even commercial) technologies but also establishes a general framework for engineering high-performance SPDs using low-technology components. This opens a new avenue for research into SPDs and related quantum optical technologies.
2 Results
In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance starting from a base (existing) SPD. Then we establish the theoretical framework governing its operation and provide analytical approximations to characterize its performance.
2.1 Design of the Enhanced Single-Photon Detector
The ESPD architecture is designed to upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
For clarity, we define the DE as the probability of generating a positive detection event given a single-photon input, and the DCR as the probability of a positive report given a vacuum input. Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with DE $\eta_{s}$ and DCR $d_{s}$. The iteration begins with an initial detector, $\text{ESPD}_{0}$, characterized by $\eta_{0}=\eta$ and $d_{0}=d$.
The design utilizes qubit logic. The schematic of the ESPD paradigm is illustrated in Fig. 1.
Figure 1: Schematic representation of the recursive ESPD architecture. The $(s+1)$-th level detector ($\text{ESPD}_{s+1}$) is constructed by integrating multiple copies of the $s$-th level detector ($\text{ESPD}_{s}$) with controlled gates. Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective measurements via basis $\{|0\rangle,|1\rangle\}$, and filters represent that the $|0\rangle$ path is aborted.
The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, proceeds as follows. To facilitate understanding, readers can consider an implementation where the degree of freedom (DOF) is photon polarization and the controlled operations are C-NOT gates.
  1. 1.
    DOF Selection: Select a manipulable DOF of the incoming photon (e.g., polarization). Let $\{|0\rangle,|1\rangle\}$ be the computational basis for this DOF. We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon. Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum. However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from it.
  2. 2.
    State Preparation: Apply a state preparation operation to the incoming signal to initialize the selected DOF to $|1\rangle$. Consequently, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum signal remains $|\emptyset\rangle$.
  3. 3.
    Controlled Operations: Implement $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates). These gates interact the signal path with $n_{s+1}$ auxiliary sources, each initialized in the state $|0\rangle$.
  4. 4.
    Auxiliary Measurement: Perform projective measurements on the auxiliary outputs in the basis $\{|0\rangle,|1\rangle\}$. Outcome paths of $|0\rangle$ are discarded, and only the $|1\rangle$ paths are preserved for detection.
  5. 5.
    Detection: Detect the signals at the end of paths, including the signal one and the auxiliary ones, using $s$-th level detectors, $\text{ESPD}_{s}$.
  6. 6.
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detections are registered among the $n_{s+1}+1$ detectors (comprising $n_{s+1}$ auxiliary detectors and one signal path detector). Otherwise, a negative report is returned. Here, $k_{s+1}$ is a pre-selected threshold integer. Here, $n_{s+1}$ should refer to the number of effective operations that survive post-selection (if the controlled gates include such), namely, the decision logic is essentially that the positive report rate on the multiple detections reaches a pre-defined threshold.
2.2 Theoretical Formulation
2.2.1 Notations
We define the parameters governing the system dynamics. As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may theoretically depend on the photon number distribution, we approximate it as a unified value for the given system. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled gates and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
We assume all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$ for convenience. Let $p$ be the transmission probability of the signal path through a single module (an optional state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Furthermore, let $Q$ represent the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement, but not including detection). Given the maturity of commercial linear optical components, we assume $Q\ll 1$ (typically $Q<10^{-3}$). Finally, let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. For simplicity, we approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ operations.
2.2.2 Assumptions
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. We further assume $d_{s}\ll P\eta_{s}$. This condition is trivial, for if it were violated, the signal contribution from the controlled operations would be indistinguishable from device noise, rendering the operations practically insignificant. Additionally, the transmission rate is assumed to be $p\approx 1$ (see Section 3.1). Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
2.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to the $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. These are expressed as:
$P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
$Q_{s+1}=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$ (2)
Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final gate, respectively, then:
$P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\ Q^{\prime}_{s+1}=d_{s}$ (3)
We now derive the DE and DCR for the $(s+1)$-th level.
For the DE calculation, we assume a non-vacuum input and calculate the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections. If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
$P_{s+1,k_{s+1},i}$ (4)
$=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
$+Q^{\prime}_{s+1}\sum_{j_{1}+j_{2}=k_{s+1}-1}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
where the first and second terms correspond to scenarios where the signal path detector reports negative or positive, respectively.
If the photon survives all $C(F)$ operations, the probability is:
$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
$\qquad\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
Combining these, the DE of $\text{ESPD}_{s+1}$ is:
$\eta_{s+1}=p^{n_{s+1}}P_{s+1,k_{s+1}}+\sum_{i=1}^{n_{s+1}}p^{i-1}(1-p)P_{s+1,k_{s+1},i}$ (6)
For the DCR calculation, we assume a vacuum input. The probability of a false positive is:
$d_{s+1}=(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
$\qquad\qquad\qquad\qquad\qquad\qquad\quad+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
Collectively, the DE and DCR evolution is governed by the non-linear time-variant dynamic system:
$(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\ (\eta_{0},d_{0})=(\eta,d)$ (8)
where the mapping $G$ is defined by Eq. (1) to (7).
2.3 Analytical Approximation and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, employing the approximations in Eq.(1) to (3) and the assumptions from Section 2.2.2, we can derive simplified bounds to elucidate the system’s behavior.
Consider the DCR evolution. Eq. (7) can be reformulated as:
$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}$ (9)
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
Proposition 1.
Define the function
$f(x)=a\binom{n}{k-1}x^{k-1}(1-x)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}x^{j}(1-x)^{n-j}$ (10)
Then $f(x)$ is monotonically increasing for $x\in[0,\frac{k-1}{n}]$.
Proof: For $x\in[0,\frac{k-1}{n}]$:
$\frac{d}{dx}f(x)$ (11)
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$
$\qquad\qquad\qquad\quad+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}(k-1-nx)+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}(j-nx)]$
$\geq$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}](k-1-nx)\geq 0$
$\blacksquare$
Applying the upper bound from Eq. (2), substituting Eq. (3) to Eq. (9), and assuming $k-1\geq nx$, we obtain:
$d_{s+1}\leq d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}$ (12)
$\qquad\qquad\qquad\qquad\qquad\qquad\quad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$
which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate $n_{s+1},k_{s+1}$, we assume $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
$d_{s+1}$ $\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}$ (13)
$=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$
This provides an estimation for the DCR at level $s+1$.
Similarly, the DE can be lower-bounded by:
$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (14)
$\qquad\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
which is independent of $d_{s}$. Therefore, solving the inequality
$p^{n}(Px)^{k-1}[x\binom{n}{k-1}(1-Px)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}(Px)^{j-k+1}(1-Px)^{n-j}]-x>0$ (15)
determines the condition under which the DE is enhanced in the subsequent level. If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
2.4 Performance Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, we present numerical simulations to validate the ESPD paradigm, demonstrating that substantial performance gains are achievable even with moderate initial hardware specifications. For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $95\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
Simulation Parameters: We evaluated the ESPD scheme using two baseline detectors: one representing older technology with $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], and another representing typical InGaAs/InP performance with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. The transmission probability is set to $p=0.99$, and the auxiliary path error rate $Q=0.2\%$ (see Section 3.1). We considered two controlled gate fidelity scenarios: $P=0.81$, representing standard linear optical C-NOT gates [38], and P=0.95P=0.95111Corresponding to about 0.22 dB, which is very likely to be achieved commercially in the short term, since it is confident to achieve by the commercial platform Q.ANT in 2026., representing high-fidelity implementations [39, 40]. Also, a further implementation with $P$ reduced to 0.4 is reported in Section 3.1. Various combinations of the configuration parameters $n_{s}$ and $k_{s}$ are evaluated.
Performance Outcomes: Table 2 to 5 and Fig. 2 to 5 summarize the achievable performance metrics across varying configurations and baseline architectures. The results indicate that the ESPD scheme significantly enhances the capabilities of conventional SPDs. Notably, with only three levels of enhancement, the detectors achieve performance metrics (DE $95\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
3 Discussion
We now discuss the device requirements, potential applications, as well as current limitations and avenues for future work.
Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and high controlled gate fidelity, $P=0.99$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,2) $(88.7\%,\ 1.2\times 10^{-3})$ (6,3) $(86.2\%,\ 4.5\times 10^{-5})$ (7,3) $(90.9\%,\ 7.2\times 10^{-5})$
2 (4,2) $(98.6\%,\ 6.5\times 10^{-5})$ (6,3) $(97.4\%,\ 1.1\times 10^{-8})$ (7,3) $(97.7\%,\ 2.4\times 10^{-7})$
3 (4,2) $(99.0\%,\ 2.5\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
4 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
5 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
6 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
7 (4,2) $(99.0\%\ ,2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
8 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
Para 4 Para 5 Para 6
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (8,4) $(84.3\%,\ 1.8\times 10^{-6})$ (8,2) $(97.3\%,\ 4.2\times 10^{-3})$ (8,2) $(97.3\%,\ 4.2\times 10^{-3})$
2 (8,4) $(96.2\%,5.7\times 10^{-10})$ (8,4) $(96.9\%,\ 1.5\times 10^{-7})$ (8,3) $(97.9\%,\ 1.7\times 10^{-5})$
3 (8,4) $(96.8\%,9.5\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,\ 1.1\times 10^{-9})$
4 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
5 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
6 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
7 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
8 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 2: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.99$.
Table 3: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.81$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,2) $(80.2\%,\ 1.2\times 10^{-3})$ (6,3) $(74.2\%,\ 4.5\times 10^{-5})$ (7,3) $(81.6\%,\ 7.2\times 10^{-5})$
2 (4,2) $(94.5\%,\ 5.9\times 10^{-5})$ (6,3) $(89.1\%,\ 7.3\times 10^{-8})$ (7,3) $(95.2\%,\ 1.8\times 10^{-7})$
3 (4,2) $(97.9\%,\ 2.3\times 10^{-5})$ (6,3) $(95.7\%,\ 1.1\times 10^{-7})$ (7,3) $(97.0\%,\ 2.4\times 10^{-7})$
4 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(96.8\%,\ 1.4\times 10^{-7})$ (7,3) $(97.2\%,\ 2.5\times 10^{-7})$
5 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(96.9\%,\ 1.4\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
6 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
7 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
8 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
Para 4 Para 5 Para 6
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (8,4) $(69.6\%,\ 1.8\times 10^{-6})$ (8,2) $(95.2\%,\ 4.2\times 10^{-3})$ (8,2) $(95.2\%,\ 4.2\times 10^{-3})$
2 (8,4) $(82.9\%,2.6\times 10^{-10})$ (8,4) $(95.6\%,\ 1.4\times 10^{-7})$ (8,3) $(97.2\%,\ 1.6\times 10^{-5})$
3 (8,4) $(92.5\%,5.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.8\%,\ 1.0\times 10^{-9})$
4 (8,4) $(95.2\%,8.2\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.7\%,9.4\times 10^{-10})$
5 (8,4) $(95.6\%,9.1\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
6 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
7 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
8 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 3: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.81$.
Table 4: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$, and high controlled gate fidelity, $P=0.99$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,1) $(78.6\%,\ 2.2\times 10^{-3})$ (6,1) $(87.6\%,\ 3.3\times 10^{-3})$ (8,1) $(92.3\%,\ 4.4\times 10^{-3})$
2 (4,2) $(97.5\%,\ 1.2\times 10^{-4})$ (6,2) $(98.7\%,\ 4.7\times 10^{-4})$ (8,4) $(96.7\%,\ 1.6\times 10^{-7})$
3 (4,2) $(98.9\%,\ 2.7\times 10^{-5})$ (6,3) $(98.0\%,\ 3.7\times 10^{-7})$ (8,4) $(96.9\%,\ 9.7\times 10^{-10})$
4 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
5 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4)) $(96.9\%,\ 9.8\times 10^{-10})$
6 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
7 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
8 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 4: Performance scaling for initial parameters $\eta_{0}=27.5\%$, $d_{0}=1.0\times 10^{-6}$, with $P=0.99$.
Table 5: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$, and moderate controlled gate fidelity, $P=0.81$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (4,1) $(72.4\%,\ 2.2\times 10^{-3})$ (6,1) $(82.3\%,\ 3.3\times 10^{-3})$ (8,1) $(88.1\%,\ 4.4\times 10^{-3})$
2 (4,2) $(90.7\%,\ 1.1\times 10^{-4})$ (6,2) $(97.7\%,\ 4.6\times 10^{-4})$ (8,4) $(94.4\%,\ 1.5\times 10^{-7})$
3 (4,2) $(97.3\%,\ 2.3\times 10^{-5})$ (6,3) $(97.0\%,\ 3.2\times 10^{-7})$ (8,4) $(95.5\%,\ 8.8\times 10^{-10})$
4 (4,2) $(98.2\%,\ 2.3\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
5 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4)) $(95.6\%,\ 9.3\times 10^{-10})$
6 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
7 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
8 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 5: Performance scaling for initial parameters $\eta_{0}=27.5\%$, $d_{0}=1.0\times 10^{-6}$, with $P=0.81$.
Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.86 to 0.98. Here, $P=0.99$ and experiments are run for $(n,k)=(5,2)$ on each level.
pp=0.86 pp=0.89 pp=0.92 pp=0.95 pp=0.98
Level (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$))
0 $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$
1 $(70.5\%,1.8\times 10^{-3})$ $(75.3\%,1.8\times 10^{-3})$ $(80.4\%,1.8\times 10^{-3})$ $(85.8\%,1.8\times 10^{-3})$ $(91.4\%,1.8\times 10^{-3})$
2 $(76.3\%,1.3\times 10^{-4})$ $(82.6\%,1.3\times 10^{-4})$ $(88.4\%,1.4\times 10^{-4})$ $(93.4\%,1.5\times 10^{-4})$ $(97.6\%,1.6\times 10^{-4})$
3 $(78.6\%,2.8\times 10^{-5})$ $(84.8\%,3.3\times 10^{-5})$ $(90.0\%,3.8\times 10^{-5})$ $(94.3\%,4.2\times 10^{-5})$ $(97.9\%,4.6\times 10^{-5})$
4 $(79.5\%,2.6\times 10^{-5})$ $(85.4\%,3.0\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
5 $(79.9\%,2.6\times 10^{-5})$ $(85.6\%,3.0\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
6 $(80.0\%,2.6\times 10^{-5})$ $(85.6\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
7 $(80.0\%,2.7\times 10^{-5})$ $(85.6\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
8 $(80.0\%,2.7\times 10^{-5})$ $(85.7\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
(a) Detection efficiency $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
(b) Dark count rate $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
Figure 6: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.99$.
Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under a low controlled gate fidelity, $P=0.40$.
Para 1 Para 2 Para 3
Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
[…]
1 (6,2) $(64.0\%,\ 2.4\times 10^{-3})$ (7,2) $(69.8\%,\ 3.3\times 10^{-3})$ (8,2) $(74.5\%,\ 4.2\times 10^{-3})$
2 (6,2) $(68.0\%,\ 2.6\times 10^{-4})$ (7,2) $(78.5\%,\ 5.5\times 10^{-4})$ (8,2) $(85.5\%,\ 1.1\times 10^{-3})$
3 (6,2) $(71.5\%,\ 4.2\times 10^{-5})$ (7,2) (84.4%, $4\%,\ 1.0\times 10^{-4}) (8$,2) $(90.6\%,\ 2.3\times 10^{-4})$
4 (6,2) $(74.7\%,\ 3.3\times 10^{-5})$ (7,2) (87.7%, $7\%,\ 6.8\times 10^{-5}) (8$,2) $(92.3\%,\ 1.2\times 10^{-4})$
5 (6,2) $(77.4\%,\ 3.5\times 10^{-5})$ (7,2) $(89.4\%,\ 7.0\times 10^{-5})$ (8,2) $(92.9\%,\ 1.1\times 10^{-4})$
6 (6,2) $(79.5\%,\ 3.8\times 10^{-5})$ (7,2) $(90.1\%,\ 7.3\times 10^{-5})$ (8,2) $(93.0\%,\ 1.1\times 10^{-4})$
7 (6,2) $(81.0\%,\ 4.0\times 10^{-5})$ (7,2) $(90.4\%,\ 7.4\times 10^{-5})$ (8,2) $(93.1\%,\ 1.1\times 10^{-4})$
8 (6,2) $(82.1\%,\ 4.2\times 10^{-5})$ (7,2) $(90.5\%,\ 7.5\times 10^{-5})$ (8,2) $(93.1\%,\ 1.1\times 10^{-4})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 7: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
3.1 Device Requirements
The flexibility of the ESPD paradigm, which allows for various DOF to be employed, enables the selection of an easily manipulated parameter. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates. Specifically, $Q$ can be suppressed to below the $10^{-4}$ level using commercial platforms[41, 42], thereby guaranteeing condition $Q\ll 1$.
The primary technological requirements of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$. In practice, the transmission rate $p$ is often approximated as $p\approx 1$, representing a high transmission efficiency [43, 44, 45, 46, 47, 48, 49, 42]. However, since $p$ decays exponentially as the number of stages $n$ increases, experimental characterization of $p$ remains crucial. Table 6 and Fig.6 demonstrate the performance when $p$ is degraded, showing that the ESPD framework remains valid even if $p$ drops to $0.86$, albeit with reduced performance compared to higher values.
Attention now shifts to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities, exceeding $99\%$ [39, 40] using modern technology and reaching $81\%$ via linear-optical methods [38], suggesting that high values of $P$ are readily attainable. Our numerical results in Section 2.4 confirm that the ESPD paradigm does not rely on an excessively demanding value for $P$, showing similar substantial performance gains for $P=0.99$ and $P=0.81$. Furthermore, Table 7 and Fig.7 illustrate the robustness of the scheme, confirming its efficacy even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This demonstrated robustness confirms that the technological demands of the ESPD method are fully compatible with existing quantum technologies.
It is also important to note that the implementation of the ESPD scheme does not require perfect auxiliary photon sources. The parameter $P$ relies solely on the detection of the state $|1\rangle$ from the auxiliary path when the control signal input is non-vacuum. This condition holds true regardless of the exact number of photons in the auxiliary path or the presence of unwanted $|0\rangle$ measurement outcomes, which are inherently filtered out [^2] This justifies the necessity of the dedicated filter block within the design.. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required final $|1\rangle$, although a rigorous theoretical formulation of this potential advantage is reserved for future work.
Finally, the requirements for the initial single-photon detector (SPD) used for initialization are minimal. As demonstrated in Section 2.4, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval. The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point. Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$. One or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient. Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable starting points. In summary, the ESPD scheme only requires a normal, low-performance SPD for initialization, which is easily achieved. For instance, a device dating back two decades [28] is sufficient.
3.2 Applications
To illustrate the practical impact of the ESPD scheme, we consider its application in QKD. For a QKD protocol having a secure threshold $e_{th}$, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
$\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude, it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor. Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency. This allows for significant QKD performance improvements without an additional efficiency cost. Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
3.3 Limitation and Future Work
Despite the significance of the ESPD framework, some limitations remain that are left for future investigations.
3.3.1 Theoretical Completeness
The theoretical framework established in Section 2.2 remains an area for deeper exploration. While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest. Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required. Given the inherent mathematical difficulty of such a non-linear and time-variant system, these specialized studies are reserved for the future.
3.3.2 Practical Cost and Optimization
Although the parallelized nature of the ESPD’s detection stages theoretically preserves detection efficiency, the requirement for multiple detection events introduces a practical cost and operational burden, even without increasing the fundamental technological barriers. An $L$-level ESPD, with $n$ auxiliary signals per level, requires a total of $(n+1)^{L}$ base detections. While effective performance is typically achieved with only a few levels, this inherent multiplicative complexity increases the practical cost, which may limit immediate commercial scalability. As shown in Section 2.4, the optimal choice of $n$ and $k$ is crucial as it dictates the convergence speed and the resulting DE-DCR trade-off. Enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and larger $n$, and enlarging $n$ increases detection cost. Therefore, designing the ESPD with optimal parameters constitutes a significant multi-objective optimization problem for the future.
This optimization problem also encompasses the optimal circuit design. To mitigate error propagation, minimizing the maximal circuit length is essential. Optimal designs could involve leveraging a binary-tree structure and defining the positive report event via a global, rather than the current level-by-level local, decision criterion, which expands the tree structure used in Section 2.1 with a local decision criterion. However, both problems require further and specific discussions: the general binary-tree structure results in non-independent distributions that are difficult to analyze, while the global decision criterion design must account for both device requirements and implementation convenience.
3.3.3 Physical Integrations and Experiments
Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, the physical implementation of the ESPD scheme requires dedicated experimental validation. Furthermore, the seamless integration of controlled gates into various quantum optical tasks has not yet been fully investigated. Thus, both the physical realization of the ESPD scheme and its application to a wide range of tasks are key directions for future work. Reassuringly, as established in Section 3.1, no fundamental technological barriers prevent its current achievement.
4 Conclusion
In this paper, we have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for transforming low-performance, conventional SPDs into devices with performance comparable to or exceeding state-of-the-art superconducting SPDs. The ESPD scheme is characterized by significantly reduced technology requirements and immediate physical realizability using current quantum components. Our numerical analysis demonstrates that a legacy non-superconducting SPD, with DE $\approx 59\%$ and DCR $\approx 10^{-2}$, can be enhanced to one achieving DE $95\%$ and DCR $10^{-9}$. This dramatic boost circumvents the need for cryogenic superconducting materials, enabling the deployment of high-performance detection capabilities outside specialized laboratories. Furthermore, this enhancement allows the minimal tolerable channel transmission rate in quantum communication to be reduced by multiple orders of magnitude. The ESPD framework thus paves a clear, scalable path toward the commercialization and widespread implementation of high-performance quantum optical tasks.
Declarations
The author declares no financial or non-financial competing interests.
Author Contributions
Hao Shu (the unique author): study design, investigation, experiment, analysis, manuscript preparation.
Data and Code Availability
All data/codes generated or used during this study are included in this published article or can be found at https://github.com/Hao-B-Shu/ESPD.

### v2

@@
-Enhanced Single-Photon Detector: Achieving Superconducting-Level Performance with Conventional Quantum Technology
+Enhanced Single-Photon Detector: A framework for Superconducting-Level Performance without cryogenic cooling
 Abstract
-High-performance single-photon detectors (SPDs) are indispensable components for quantum optical tasks.
-However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and challenging operational conditions (e.g., cryogenics), which hinder scalable commercial deployment.
-To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of low-technology SPDs.
-Utilizing only standard quantum optical components, the ESPD scheme transforms a legacy non-superconducting SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) $10^{-2}$, into a device with superior performance metrics, achieving DE higher than $95\%$ and DCR below $10^{-9}$.
-This level of performance is comparable to or surpasses recently designed superconducting SPDs, allowing the minimal tolerable channel transmission rate for Quantum Key Distribution (QKD) protocols to be reduced by several orders of magnitude.
-Furthermore, the scheme’s device requirements are moderate, relying on readily available current technology, which ensures near-term experimental feasibility.
-The ESPD framework thus provides a clear, scalable path toward the large-scale deployment of high-performance SPDs and the commercialization of quantum communication technologies.
+High-performance single-photon detectors (SPDs) are indispensable components for a wide range of quantum optical applications.
+However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and necessitates challenging operational conditions, such as cryogenics, thereby hindering scalable implementation.
+To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of conventional room-temperature SPDs.
+Relying entirely on non-superconducting components and eliminating cryogenic cooling requirements, the ESPD scheme can upgrade a legacy SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) of $10^{-2}$, to a device with superior performance metrics, achieving DE exceeding $93\%$ and DCR below $10^{-9}$.
+This performance rivals or even surpasses that of state-of-the-art superconducting SPDs, allowing the minimal tolerable channel transmission rate for quantum key distribution (QKD) protocols to be reduced by several orders of magnitude.
+Although the architecture requires substantial integration effort, the scheme can provide superconducting-level performance without cryogenic cooling, offering a clear path toward the widespread deployment of high-performance SPDs as well as related quantum technologies in infrastructure-constrained environments.
 keywords:
-Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
+Room-Temperature Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
 1 Introduction
-Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurement is indispensable for information readout across virtually all applications.
+Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurements are indispensable for information readout across virtually all applications.
 The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$.
-The former influences the efficiency of nearly all quantum operations, while the latter constitutes the ultimate limitation in high-attenuation scenarios.
-Regarding DE, numerous quantum experiments rely on SPD signals to retrieve results.
+The former affects the efficiency of nearly all quantum operations, while the latter imposes the ultimate limitation in high-attenuation scenarios.
+In terms of DE, numerous quantum experiments rely on SPD signals to retrieve results.
 In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3].
 In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12].
-In heralded single-photon sources, where entangled pairs are generated and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16].
-Regarding DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD).
-The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, which is significantly lower than 50% [17, 18, 19, 20].
+In heralded single-photon sources, where entangled pairs are generated, and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16].
+With respect to DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD).
+The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, significantly lower than 50% [17, 18, 19, 20].
 However, communication channels induce exponential photon attenuation [21, 22, 23, 24].
@@
 Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
-Given their critical importance, the development of SPDs has been a continuous research focus since the inception of quantum technologies.
+Driven by their critical importance, SPD has remained a continuous research focus since the inception of quantum technologies.
 While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR.
-For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP detector presented in 2017 operates at 223 K with a DE of 27.5% and a DCR of $10^{-6}$ [29].
-These specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$.
+For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with DE about 59% and DCR of $10^{-2}$ [28], while an InGaAs/InP SPD presented in 2017 operates at 223 K with DE about 27.5% and DCR of $10^{-6}$ [29].
+Such specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$.
 To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33].
-These detectors achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude.
-This capability enables the implementation of demanding quantum tasks, such as long-distance QKD distance [19, 34, 35, 36].
+These SPDs achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude.
+This capability enables the implementation of demanding quantum tasks, such as long-distance QKD [19, 34, 35, 36].
 A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
@@
 However, as illustrated in the table, the reliance on superconductors necessitates stringent operating conditions, specifically with extremely low temperatures.
-Furthermore, these systems often entail high technical complexity and cost, hindering scalable commercial deployment.
-To address these limitations, we propose a design for SPDs that meets advanced performance requirements, and even potentially outperforms superconducting devices, while relying on accessible, non-superconducting technology suitable for room-temperature operation.
-The proposed method, termed the Enhanced SPD (ESPD) scheme, iteratively upgrades the performance of an existing SPD level by level.
-As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] to one achieving a DE exceeding 95% and a DCR below $10^{-9}$, thereby rivaling state-of-the-art devices.
-Our method not only provides a high-performance SPD design realizable with current (even commercial) technologies but also establishes a general framework for engineering high-performance SPDs using low-technology components.
-This opens a new avenue for research into SPDs and related quantum optical technologies.
-2 Results
-In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance starting from a base (existing) SPD.
-Then we establish the theoretical framework governing its operation and provide analytical approximations to characterize its performance.
+When these demanding conditions are relaxed, the SPD performance degrades significantly.
+This requirement introduces significant infrastructural requirements and operational complexity, fundamentally hindering the scalability and widespread implementation of high-performance quantum technology in field-deployable and environmentally constrained situations.
+To overcome these limitations, we propose the enhanced SPD (ESPD) architecture that can meet advanced performance requirements, potentially outperforming superconducting ones, while relying solely on non-superconducting components suitable for room-temperature operation.
+The proposed method iteratively upgrades the performance of an existing conventional SPD level by level, ultimately to a high-performance one.
+As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] with DE about 59% and DCR of $10^{-2}$ to one achieving DE exceeding 93% and DCR below $10^{-9}$, thereby can rival the state-of-the-art devices and reduce the minimal tolerable channel transmission rate in QKD by several orders of magnitude.
+Despite the substantial engineering integration effort required for its large-scale architecture, the ESPD method offers two key advancements: It not only provides a high-performance SPD design approach without cryogenic components, enabling a practical route toward accurate room-temperature photonic detection, but also establishes a paradigm shift for quantum information processing by leveraging multi-copy analysis and iterative enhancement to boost performance using conventional components.
+Consequently, this work opens a new avenue for the scalable deployment of high-performance SPDs and advances broader quantum optical technologies in infrastructure-constrained environments.
+2 Result
+In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance initialized from a base (existing) SPD.
+We establish the theoretical framework governing its operation, provide analytical approximations to characterize its performance, and state its applications in QKD.
 2.1 Design of the Enhanced Single-Photon Detector
-The ESPD architecture is designed to upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
-For clarity, we define the DE as the probability of generating a positive detection event given a single-photon input, and the DCR as the probability of a positive report given a vacuum input.
+The ESPD architecture is designed to recursively upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
+For clarity, we clarify that DE is the probability of generating a positive detection event given a non-vacuum (single-photon) input, and the DCR is the probability of a positive report given a vacuum input.
 Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with DE $\eta_{s}$ and DCR $d_{s}$.
-The iteration begins with an initial detector, $\text{ESPD}_{0}$, characterized by $\eta_{0}=\eta$ and $d_{0}=d$.
+The iteration begins with an initial existing SPD, $\text{ESPD}_{0}$, characterized by initial value $\eta_{0}=\eta$ and $d_{0}=d$.
 The design utilizes qubit logic. The schematic of the ESPD paradigm is illustrated in Fig. 1.
@@
     DOF Selection: Select a manipulable DOF of the incoming photon (e.g., polarization).
-Let $\{|0\rangle,|1\rangle\}$ be the computational basis for this DOF.
-We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon.
-Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum.
-However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from it.
+Let $\{|0\rangle,|1\rangle\}$ denote the computational basis for this DOF.
+We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon [^1] Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum, which degenerates to the qubit case.
+However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from photon existence..
   2. 2.
@@
 Otherwise, a negative report is returned.
-Here, $k_{s+1}$ is a pre-selected threshold integer.
-Here, $n_{s+1}$ should refer to the number of effective operations that survive post-selection (if the controlled gates include such), namely, the decision logic is essentially that the positive report rate on the multiple detections reaches a pre-defined threshold.
-2.2 Theoretical Formulation
+Here, $k_{s+1}$ is a pre-selected threshold integer, and $n_{s+1}$ refers to the number of effective operations [^2] For example, if the controlled gates are deterministic, then $n_{k+1}$ is exactly the number of used operations, while if the controlled gates are probabilistic, then $n_{k+1}$ refers to the successful ones after post-selection.
+Therefore, the decision logic is essentially that the positive report rate (among all effective reports) on the multiple detections reaches a pre-defined threshold..
+2.2 Theoretical Formulations
 2.2.1 Notations
@@
 As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively.
-Although $\eta_{s}$ may theoretically depend on the photon number distribution, we approximate it as a unified value for the given system.
-The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled gates and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
-We assume all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$ for convenience.
-Let $p$ be the transmission probability of the signal path through a single module (an optional state preparation followed by a $C(F)$).
-Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input, accounting for loss but irrespective of decoherence (since the state can be re-prepared).
-Furthermore, let $Q$ represent the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement, but not including detection).
-Given the maturity of commercial linear optical components, we assume $Q\ll 1$ (typically $Q<10^{-3}$).
-Finally, let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively.
-For simplicity, we approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ operations.
+Although $\eta_{s}$ may depend on the photon number distribution, we approximate it as a unified value for the given system for simplicity.
+The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
+For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$.
+Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$).
+Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared).
+Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively.
+Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection).
+We approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ module operations.
+It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module.
+These parameters are defined at the component level and remain invariant with respect to the number of cascaded modules.
+The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
 2.2.2 Assumptions
@@
 Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds.
-We further assume $d_{s}\ll P\eta_{s}$.
-This condition is trivial, for if it were violated, the signal contribution from the controlled operations would be indistinguishable from device noise, rendering the operations practically insignificant.
-Additionally, the transmission rate is assumed to be $p\approx 1$ (see Section 3.1).
+We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant.
+Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section 3.1).
+Moreover, for analytical simplicity, we assume the controlled gates are implemented in a feedforward controlled configuration.
+This ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$.
+However, we emphasize that the generalized ESPD design remains valid for post-selection controlled gates (for a brief discussion, see Section 3.1.2).
 Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
 2.2.3 Formulations
-Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to the $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector.
-These are expressed as:
+Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector.
+They can be expressed as:
 $P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
 $Q_{s+1}=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$ (2)
-Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final gate, respectively, then:
+Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final controlled gate, respectively, then:
 $P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\ Q^{\prime}_{s+1}=d_{s}$ (3)
@@
 $+Q^{\prime}_{s+1}\sum_{j_{1}+j_{2}=k_{s+1}-1}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
-where the first and second terms correspond to scenarios where the signal path detector reports negative or positive, respectively.
-If the photon survives all $C(F)$ operations, the probability is:
+where the first and second terms correspond to scenarios that the signal path detector reports negative and positive, respectively.
+If the photon survives all $C(F)$ operations, the probability of a final positive report is:
 $P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
@@
 $(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\ (\eta_{0},d_{0})=(\eta,d)$ (8)
-where the mapping $G$ is defined by Eq. (1) to (7).
-2.3 Analytical Approximation and Stability Analysis
+where the map $G_{s+1}$ is defined by Eq. (1) to (7).
+2.3 Analytical Approximations and Stability Analysis
 While Eq.
@@
 2.4 Performance Simulations
-While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
+While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
 In this section, we present numerical simulations to validate the ESPD paradigm, demonstrating that substantial performance gains are achievable even with moderate initial hardware specifications.
-For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $95\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
-Simulation Parameters: We evaluated the ESPD scheme using two baseline detectors: one representing older technology with $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], and another representing typical InGaAs/InP performance with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
-The transmission probability is set to $p=0.99$, and the auxiliary path error rate $Q=0.2\%$ (see Section 3.1).
-We considered two controlled gate fidelity scenarios: $P=0.81$, representing standard linear optical C-NOT gates [38], and P=0.95P=0.95111Corresponding to about 0.22 dB, which is very likely to be achieved commercially in the short term, since it is confident to achieve by the commercial platform Q.ANT in 2026., representing high-fidelity implementations [39, 40].
-Also, a further implementation with $P$ reduced to 0.4 is reported in Section 3.1.
-Various combinations of the configuration parameters $n_{s}$ and $k_{s}$ are evaluated.
-Performance Outcomes: Table 2 to 5 and Fig.
-2 to 5 summarize the achievable performance metrics across varying configurations and baseline architectures.
-The results indicate that the ESPD scheme significantly enhances the capabilities of conventional SPDs.
-Notably, with only three levels of enhancement, the detectors achieve performance metrics (DE $95\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
-3 Discussion
-We now discuss the device requirements, potential applications, as well as current limitations and avenues for future work.
-Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and high controlled gate fidelity, $P=0.99$.
+For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
+Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs.
+The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
+Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation (see Section 3.1).
+We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section 3.1).
+Performance Outcomes: Achievable performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig.
+2 and 3.
+The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs.
+Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 2: Performance for initial parameters $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
+Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
 Para 1 Para 2 Para 3
@@
 […]
-1 (4,2) $(88.7\%,\ 1.2\times 10^{-3})$ (6,3) $(86.2\%,\ 4.5\times 10^{-5})$ (7,3) $(90.9\%,\ 7.2\times 10^{-5})$
-2 (4,2) $(98.6\%,\ 6.5\times 10^{-5})$ (6,3) $(97.4\%,\ 1.1\times 10^{-8})$ (7,3) $(97.7\%,\ 2.4\times 10^{-7})$
-3 (4,2) $(99.0\%,\ 2.5\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
-4 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
-5 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
-6 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
-7 (4,2) $(99.0\%\ ,2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
-8 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
-Para 4 Para 5 Para 6
-Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
-[…]
-1 (8,4) $(84.3\%,\ 1.8\times 10^{-6})$ (8,2) $(97.3\%,\ 4.2\times 10^{-3})$ (8,2) $(97.3\%,\ 4.2\times 10^{-3})$
-2 (8,4) $(96.2\%,5.7\times 10^{-10})$ (8,4) $(96.9\%,\ 1.5\times 10^{-7})$ (8,3) $(97.9\%,\ 1.7\times 10^{-5})$
-3 (8,4) $(96.8\%,9.5\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,\ 1.1\times 10^{-9})$
-4 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
-5 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
-6 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
-7 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
-8 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
-(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 2: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.99$.
-Table 3: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.81$.
+1 (4,1) $(97.4\%,\ 5.3\times 10^{-2})$ (6,1) $(98.4\%,\ 7.5\times 10^{-2})$ (8,1) $(98.6\%,\ 9.5\times 10^{-2})$
+2 (4,2) $(98.2\%,\ 2.7\times 10^{-2})$ (6,3) $(96.6\%,\ 1.2\times 10^{-2})$ (8,4) $(95.1\%,\ 7.4\times 10^{-3})$
+3 (4,2) $(98.0\%,\ 7.7\times 10^{-3})$ (6,3) $(95.8\%,\ 8.9\times 10^{-5})$ (8,4) $(93.6\%,\ 8.1\times 10^{-7})$
+4 (4,2) $(97.9\%,\ 8.4\times 10^{-4})$ (6,3) $(95.6\%,\ 1.7\times 10^{-7})$ (8,4) $(93.4\%,\ 8.6\times 10^{-10})$
+5 (4,2) $(97.8\%,\ 5.6\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+6 (4,2) $(97.8\%,\ 2.5\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+7 (4,2) $(97.8\%\ ,2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 3: Performance for initial parameters $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
+Table 3: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
 Para 1 Para 2 Para 3
@@
 […]
-1 (4,2) $(80.2\%,\ 1.2\times 10^{-3})$ (6,3) $(74.2\%,\ 4.5\times 10^{-5})$ (7,3) $(81.6\%,\ 7.2\times 10^{-5})$
-2 (4,2) $(94.5\%,\ 5.9\times 10^{-5})$ (6,3) $(89.1\%,\ 7.3\times 10^{-8})$ (7,3) $(95.2\%,\ 1.8\times 10^{-7})$
-3 (4,2) $(97.9\%,\ 2.3\times 10^{-5})$ (6,3) $(95.7\%,\ 1.1\times 10^{-7})$ (7,3) $(97.0\%,\ 2.4\times 10^{-7})$
-4 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(96.8\%,\ 1.4\times 10^{-7})$ (7,3) $(97.2\%,\ 2.5\times 10^{-7})$
-5 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(96.9\%,\ 1.4\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
-6 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
-7 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
-8 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
-Para 4 Para 5 Para 6
-Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
-[…]
-1 (8,4) $(69.6\%,\ 1.8\times 10^{-6})$ (8,2) $(95.2\%,\ 4.2\times 10^{-3})$ (8,2) $(95.2\%,\ 4.2\times 10^{-3})$
-2 (8,4) $(82.9\%,2.6\times 10^{-10})$ (8,4) $(95.6\%,\ 1.4\times 10^{-7})$ (8,3) $(97.2\%,\ 1.6\times 10^{-5})$
-3 (8,4) $(92.5\%,5.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.8\%,\ 1.0\times 10^{-9})$
-4 (8,4) $(95.2\%,8.2\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.7\%,9.4\times 10^{-10})$
-5 (8,4) $(95.6\%,9.1\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
-6 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
-7 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
-8 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
-(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 3: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.81$.
-Table 4: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$, and high controlled gate fidelity, $P=0.99$.
+1 (4,1) $(76.9\%,\ 2.2\times 10^{-3})$ (6,1) (85.5%, $5\%,\ 3.4\times 10^{-3}) (8$,1) $(90.0\%,\ 4.5\times 10^{-3})$
+2 (4,2) $(95.3\%,\ 1.2\times 10^{-4})$ (6,2) $(94.8\%,\ 3.8\times 10^{-4})$ (8,4) $(93.1\%,\ 1.7\times 10^{-7})$
+3 (4,2) $(97.7\%,\ 2.6\times 10^{-5})$ (6,3) $(95.5\%,\ 1.4\times 10^{-7})$ (8,4) $(93.3\%,\ 8.4\times 10^{-10})$
+4 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.4\times 10^{-10})$
+5 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+6 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+7 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+2.5 Application on QKD
+To illustrate the practical impact of the ESPD scheme, we consider its application in QKD.
+For a QKD protocol having a secure threshold $e_{th}$ and total error rate $e_{C}$ not accounting for SPDs, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
+$\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
+Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude (even compared to SNSPDs), it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor.
+Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
+Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency.
+This allows significant QKD performance improvements without additional efficiency costs.
+Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
+3 Discussion
+We now discuss the technological feasibility, as well as limitations and avenues for future work.
+3.1 Technological Feasibility and Performance under Degraded Parameters
+The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs.
+The device requirements are discussed as follows.
+3.1.1 State Preparations and Measurements
+The flexibility of the architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one.
+In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates.
+Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
+3.1.2 Controlled Gates
+The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
+A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
+It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module, and consequently, does not decrease with an increasing number of cascaded modules or cascaded levels.
+The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
+However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations.
+While the ESPD design remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded.
+Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$.
+Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate.
+Table 4 and Fig.4 demonstrate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.).
+This result indicates robustness, validating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
+(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
+(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
+Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
+Table 4: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.80 to 0.96.
+Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section 2.4.
+pp=0.80 pp=0.84 pp=0.88 pp=0.92 pp=0.96
+Level (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$))
+0 $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$
+1 $(60.8\%,1.8\times 10^{-3})$ $(66.7\%,1.8\times 10^{-3})$ $(72.9\%,1.8\times 10^{-3})$ $(79.7\%,1.8\times 10^{-3})$ $(87.0\%,1.8\times 10^{-3})$
+2 $(61.0\%,1.1\times 10^{-4})$ $(70.6\%,1.2\times 10^{-4})$ $(79.7\%,1.3\times 10^{-4})$ $(87.8\%,1.4\times 10^{-4})$ $(94.6\%,1.5\times 10^{-4})$
+3 $(61.0\%,1.8\times 10^{-5})$ $(72.6\%,2.4\times 10^{-5})$ $(82.1\%,3.1\times 10^{-5})$ $(89.5\%,3.7\times 10^{-5})$ $(95.4\%,4.3\times 10^{-5})$
+4 $(60.9\%,1.5\times 10^{-5})$ $(73.5\%,2.2\times 10^{-5})$ $(82.8\%,2.8\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
+5 $(60.8\%,1.5\times 10^{-5})$ $(74.0\%,2.2\times 10^{-5})$ $(83.1\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
+6 $(60.8\%,1.5\times 10^{-5})$ $(74.2\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
+7 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
+8 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
+We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate.
+Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths.
+Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38].
+Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42].
+However, it is also worth testing the effectiveness of the design under degraded parameters to verify its suitability for standard fabrication processes with lower technology requirements.
+Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6).
+All other settings remain consistent with those described in Section 2.4.
+The numerical results confirm the robustness of the scheme: Comparison of Table 5 to Table 2 demonstrates that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
+This demonstrated robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
+Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
 Para 1 Para 2 Para 3
@@
 […]
-1 (4,1) $(78.6\%,\ 2.2\times 10^{-3})$ (6,1) $(87.6\%,\ 3.3\times 10^{-3})$ (8,1) $(92.3\%,\ 4.4\times 10^{-3})$
-2 (4,2) $(97.5\%,\ 1.2\times 10^{-4})$ (6,2) $(98.7\%,\ 4.7\times 10^{-4})$ (8,4) $(96.7\%,\ 1.6\times 10^{-7})$
-3 (4,2) $(98.9\%,\ 2.7\times 10^{-5})$ (6,3) $(98.0\%,\ 3.7\times 10^{-7})$ (8,4) $(96.9\%,\ 9.7\times 10^{-10})$
-4 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
-5 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4)) $(96.9\%,\ 9.8\times 10^{-10})$
-6 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
-7 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
-8 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
-(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 4: Performance scaling for initial parameters $\eta_{0}=27.5\%$, $d_{0}=1.0\times 10^{-6}$, with $P=0.99$.
-Table 5: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$, and moderate controlled gate fidelity, $P=0.81$.
-Para 1 Para 2 Para 3
-Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
-[…]
-1 (4,1) $(72.4\%,\ 2.2\times 10^{-3})$ (6,1) $(82.3\%,\ 3.3\times 10^{-3})$ (8,1) $(88.1\%,\ 4.4\times 10^{-3})$
-2 (4,2) $(90.7\%,\ 1.1\times 10^{-4})$ (6,2) $(97.7\%,\ 4.6\times 10^{-4})$ (8,4) $(94.4\%,\ 1.5\times 10^{-7})$
-3 (4,2) $(97.3\%,\ 2.3\times 10^{-5})$ (6,3) $(97.0\%,\ 3.2\times 10^{-7})$ (8,4) $(95.5\%,\ 8.8\times 10^{-10})$
-4 (4,2) $(98.2\%,\ 2.3\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
-5 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4)) $(95.6\%,\ 9.3\times 10^{-10})$
-6 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
-7 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
-8 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
-(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) Dark count rate $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 5: Performance scaling for initial parameters $\eta_{0}=27.5\%$, $d_{0}=1.0\times 10^{-6}$, with $P=0.81$.
-Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.86 to 0.98. Here, $P=0.99$ and experiments are run for $(n,k)=(5,2)$ on each level.
-pp=0.86 pp=0.89 pp=0.92 pp=0.95 pp=0.98
-Level (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$))
-0 $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$
-1 $(70.5\%,1.8\times 10^{-3})$ $(75.3\%,1.8\times 10^{-3})$ $(80.4\%,1.8\times 10^{-3})$ $(85.8\%,1.8\times 10^{-3})$ $(91.4\%,1.8\times 10^{-3})$
-2 $(76.3\%,1.3\times 10^{-4})$ $(82.6\%,1.3\times 10^{-4})$ $(88.4\%,1.4\times 10^{-4})$ $(93.4\%,1.5\times 10^{-4})$ $(97.6\%,1.6\times 10^{-4})$
-3 $(78.6\%,2.8\times 10^{-5})$ $(84.8\%,3.3\times 10^{-5})$ $(90.0\%,3.8\times 10^{-5})$ $(94.3\%,4.2\times 10^{-5})$ $(97.9\%,4.6\times 10^{-5})$
-4 $(79.5\%,2.6\times 10^{-5})$ $(85.4\%,3.0\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
-5 $(79.9\%,2.6\times 10^{-5})$ $(85.6\%,3.0\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
-6 $(80.0\%,2.6\times 10^{-5})$ $(85.6\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
-7 $(80.0\%,2.7\times 10^{-5})$ $(85.6\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
-8 $(80.0\%,2.7\times 10^{-5})$ $(85.7\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
-(a) Detection efficiency $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
-(b) Dark count rate $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
-Figure 6: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.99$.
+1 (4,2) $(77.8\%,\ 1.2\times 10^{-3})$ (6,2) $(88.5\%,\ 2.4\times 10^{-3})$ (8,2) $(92.3\%,\ 4.2\times 10^{-3})$
+2 (4,2) $(91.3\%,\ 5.7\times 10^{-5})$ (6,3) $(92.4\%,\ 2.1\times 10^{-6})$ (8,4) $(91.0\%,\ 1.4\times 10^{-7})$
+3 (4,2) $(95.7\%,\ 2.2\times 10^{-5})$ (6,3) $(93.3\%,\ 1.3\times 10^{-7})$ (8,4) $(90.6\%,\ 7.6\times 10^{-10})$
+4 (4,2) $(96.4\%,\ 2.3\times 10^{-5})$ (6,3) $(93.5\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
+5 (4,2) $(96.5\%,\ 5.5\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
+6 (4,2) $(96.5\%,\ 2.4\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
+7 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
+8 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
+Figure 6: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
 Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under a low controlled gate fidelity, $P=0.40$.
@@
 […]
-1 (6,2) $(64.0\%,\ 2.4\times 10^{-3})$ (7,2) $(69.8\%,\ 3.3\times 10^{-3})$ (8,2) $(74.5\%,\ 4.2\times 10^{-3})$
-2 (6,2) $(68.0\%,\ 2.6\times 10^{-4})$ (7,2) $(78.5\%,\ 5.5\times 10^{-4})$ (8,2) $(85.5\%,\ 1.1\times 10^{-3})$
-3 (6,2) $(71.5\%,\ 4.2\times 10^{-5})$ (7,2) (84.4%, $4\%,\ 1.0\times 10^{-4}) (8$,2) $(90.6\%,\ 2.3\times 10^{-4})$
-4 (6,2) $(74.7\%,\ 3.3\times 10^{-5})$ (7,2) (87.7%, $7\%,\ 6.8\times 10^{-5}) (8$,2) $(92.3\%,\ 1.2\times 10^{-4})$
-5 (6,2) $(77.4\%,\ 3.5\times 10^{-5})$ (7,2) $(89.4\%,\ 7.0\times 10^{-5})$ (8,2) $(92.9\%,\ 1.1\times 10^{-4})$
-6 (6,2) $(79.5\%,\ 3.8\times 10^{-5})$ (7,2) $(90.1\%,\ 7.3\times 10^{-5})$ (8,2) $(93.0\%,\ 1.1\times 10^{-4})$
-7 (6,2) $(81.0\%,\ 4.0\times 10^{-5})$ (7,2) $(90.4\%,\ 7.4\times 10^{-5})$ (8,2) $(93.1\%,\ 1.1\times 10^{-4})$
-8 (6,2) $(82.1\%,\ 4.2\times 10^{-5})$ (7,2) $(90.5\%,\ 7.5\times 10^{-5})$ (8,2) $(93.1\%,\ 1.1\times 10^{-4})$
-(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) Dark count rate $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 7: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
-3.1 Device Requirements
-The flexibility of the ESPD paradigm, which allows for various DOF to be employed, enables the selection of an easily manipulated parameter.
-In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates.
-Specifically, $Q$ can be suppressed to below the $10^{-4}$ level using commercial platforms[41, 42], thereby guaranteeing condition $Q\ll 1$.
-The primary technological requirements of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
-In practice, the transmission rate $p$ is often approximated as $p\approx 1$, representing a high transmission efficiency [43, 44, 45, 46, 47, 48, 49, 42].
-However, since $p$ decays exponentially as the number of stages $n$ increases, experimental characterization of $p$ remains crucial.
-Table 6 and Fig.6 demonstrate the performance when $p$ is degraded, showing that the ESPD framework remains valid even if $p$ drops to $0.86$, albeit with reduced performance compared to higher values.
-Attention now shifts to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate.
-Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths.
-Contemporary C-NOT gates demonstrate high fidelities, exceeding $99\%$ [39, 40] using modern technology and reaching $81\%$ via linear-optical methods [38], suggesting that high values of $P$ are readily attainable.
-Our numerical results in Section 2.4 confirm that the ESPD paradigm does not rely on an excessively demanding value for $P$, showing similar substantial performance gains for $P=0.99$ and $P=0.81$.
-Furthermore, Table 7 and Fig.7 illustrate the robustness of the scheme, confirming its efficacy even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
-This demonstrated robustness confirms that the technological demands of the ESPD method are fully compatible with existing quantum technologies.
-It is also important to note that the implementation of the ESPD scheme does not require perfect auxiliary photon sources.
-The parameter $P$ relies solely on the detection of the state $|1\rangle$ from the auxiliary path when the control signal input is non-vacuum.
-This condition holds true regardless of the exact number of photons in the auxiliary path or the presence of unwanted $|0\rangle$ measurement outcomes, which are inherently filtered out [^2] This justifies the necessity of the dedicated filter block within the design..
-Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required final $|1\rangle$, although a rigorous theoretical formulation of this potential advantage is reserved for future work.
-Finally, the requirements for the initial single-photon detector (SPD) used for initialization are minimal.
-As demonstrated in Section 2.4, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval.
+1 (2,1) $(75.1\%,\ 3.2\times 10^{-2})$ (7,2) $(66.8\%,\ 3.3\times 10^{-3})$ (8,2) $(71.1\%,\ 4.2\times 10^{-3})$
+2 (6,2) $(78.2\%,\ 2.1\times 10^{-2})$ (7,2) $(72.8\%,\ 5.4\times 10^{-4})$ (8,2) $(79.6\%,\ 1.0\times 10^{-3})$
+3 (6,2) $(79.1\%,\ 9.5\times 10^{-3})$ (7,2) $(77.2\%,\ 9.0\times 10^{-4})$ (8,2) $(84.3\%,\ 2.1\times 10^{-4})$
+4 (6,2) $(78.7\%,\ 2.4\times 10^{-3})$ (7,2) $(80.1\%,\ 5.7\times 10^{-5})$ (8,2) $(86.4\%,\ 1.0\times 10^{-4})$
+5 (6,2) (77.7%, $7\%,\ 2.9\times 10^{-5}) (7$,2) $(82.0\%,\ 5.8\times 10^{-5})$ (8,2) $(87.2\%,\ 9.5\times 10^{-5})$
+6 (6,2) $(76.7\%,\ 5.4\times 10^{-5})$ (7,2) $(83.0\%,\ 6.1\times 10^{-5})$ (8,2) $(87.5\%,\ 9.5\times 10^{-5})$
+7 (6,2) $(76.0\%,\ 3.8\times 10^{-5})$ (7,2) $(83.6\%,\ 6.3\times 10^{-5})$ (8,2) $(87.6\%,\ 9.6\times 10^{-5})$
+8 (6,2) $(75.4\%,\ 3.7\times 10^{-5})$ (7,2) $(83.9\%,\ 6.4\times 10^{-5})$ (8,2) $(87.7\%,\ 9.6\times 10^{-5})$
+3.1.3 Choice of initialized SPDs
+It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation.
+As demonstrated in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval.
 The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point.
 Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$.
-One or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient.
-Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable starting points.
-In summary, the ESPD scheme only requires a normal, low-performance SPD for initialization, which is easily achieved.
-For instance, a device dating back two decades [28] is sufficient.
-3.2 Applications
-To illustrate the practical impact of the ESPD scheme, we consider its application in QKD.
-For a QKD protocol having a secure threshold $e_{th}$, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
-$\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
-Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude, it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor.
-Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
-Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency.
-This allows for significant QKD performance improvements without an additional efficiency cost.
-Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
-3.3 Limitation and Future Work
-Despite the significance of the ESPD framework, some limitations remain that are left for future investigations.
+Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient.
+Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
+In summary, the ESPD scheme only requires a normal, room-temperature SPD for initialization, which is routinely achievable. For instance, even an SPD developed two decades ago [28] is sufficient.
+3.1.4 Auxiliary Sources
+Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, we note that implementing the ESPD scheme may not require single photon sources.
+The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$.
+Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects.
+Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary sources, which will be filtered out and thereby have no effect even if the source is multiple-photon..
+However, the formulation of this, including its impact on controlled operations, needs a rigorous discussion and is reserved for future work.
+3.2 Limitations and Future Works
+Despite the significance of the ESPD framework, some limitations remain, which are left for future investigations.
 3.2.1 Theoretical Completeness
@@
 Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required.
-Given the inherent mathematical difficulty of such a non-linear and time-variant system, these specialized studies are reserved for the future.
-3.3.2 Practical Cost and Optimization
-Although the parallelized nature of the ESPD’s detection stages theoretically preserves detection efficiency, the requirement for multiple detection events introduces a practical cost and operational burden, even without increasing the fundamental technological barriers.
-An $L$-level ESPD, with $n$ auxiliary signals per level, requires a total of $(n+1)^{L}$ base detections.
-While effective performance is typically achieved with only a few levels, this inherent multiplicative complexity increases the practical cost, which may limit immediate commercial scalability.
-As shown in Section 2.4, the optimal choice of $n$ and $k$ is crucial as it dictates the convergence speed and the resulting DE-DCR trade-off.
-Enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and larger $n$, and enlarging $n$ increases detection cost.
-Therefore, designing the ESPD with optimal parameters constitutes a significant multi-objective optimization problem for the future.
-This optimization problem also encompasses the optimal circuit design.
-To mitigate error propagation, minimizing the maximal circuit length is essential.
-Optimal designs could involve leveraging a binary-tree structure and defining the positive report event via a global, rather than the current level-by-level local, decision criterion, which expands the tree structure used in Section 2.1 with a local decision criterion.
-However, both problems require further and specific discussions: the general binary-tree structure results in non-independent distributions that are difficult to analyze, while the global decision criterion design must account for both device requirements and implementation convenience.
+Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
+3.2.2 Resource Overhead vs. Environmental Simplicity
+While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels.
+Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections.
+Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD.
+For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components.
+While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
+However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits.
+Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57].
+Notably, very-large-scale integration of quantum photonics has been experimentally realized with over 2,500 components on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes.
+Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms.
+For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
+In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation.
+The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
+The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration.
+By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost.
+This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
+3.2.3 Optimization Design
+The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments.
+Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$).
+Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence.
+As shown in Section 2.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
+Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
+Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Fig.2.
+Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
+Para 4 Para 5 Para 6
+Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
+[…]
+1 (8,2) $(95.0\%,\ 4.2\times 10^{-3})$ (4,2) $(86.1\%,\ 1.2\times 10^{-3})$ (3,1) $(95.8\%,\ 4.3\times 10^{-2})$
+2 (8,4) $(93.5\%,\ 1.4\times 10^{-7})$ (8,4) $(92.6\%,\ 6.4\times 10^{-9})$ (6,4) $(93.9\%,\ 1.2\times 10^{-4})$
+3 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.3\%,\ 8.2\times 10^{-10})$ (8,4) $(93.4\%,\ 1.2\times 10^{-9})$
+4 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.3\%,\ 8.4\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+5 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+6 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+7 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+8 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+The optimization problem also encompasses the optimal circuit topology.
+To mitigate cumulative error propagation, minimizing the maximal circuit length is essential.
+As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the design: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1.
+However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
 3.2.4 Physical Integrations and Experiments
-Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, the physical implementation of the ESPD scheme requires dedicated experimental validation.
-Furthermore, the seamless integration of controlled gates into various quantum optical tasks has not yet been fully investigated.
-Thus, both the physical realization of the ESPD scheme and its application to a wide range of tasks are key directions for future work.
-Reassuringly, as established in Section 3.1, no fundamental technological barriers prevent its current achievement.
+Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, physical implementations of the ESPD scheme require dedicated experimental validations.
+Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization.
+However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components.
+Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation.
+Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
+3.2.5 Count Rate and Timing Jitter
+While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors.
+Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices).
+However, the architecture does not impose significant additional penalties: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening.
+Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices.
+Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
 4 Conclusion
-In this paper, we have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for transforming low-performance, conventional SPDs into devices with performance comparable to or exceeding state-of-the-art superconducting SPDs.
-The ESPD scheme is characterized by significantly reduced technology requirements and immediate physical realizability using current quantum components.
-Our numerical analysis demonstrates that a legacy non-superconducting SPD, with DE $\approx 59\%$ and DCR $\approx 10^{-2}$, can be enhanced to one achieving DE $95\%$ and DCR $10^{-9}$.
-This dramatic boost circumvents the need for cryogenic superconducting materials, enabling the deployment of high-performance detection capabilities outside specialized laboratories.
-Furthermore, this enhancement allows the minimal tolerable channel transmission rate in quantum communication to be reduced by multiple orders of magnitude.
-The ESPD framework thus paves a clear, scalable path toward the commercialization and widespread implementation of high-performance quantum optical tasks.
+We have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for upgrading low-performance SPDs into devices whose capabilities can match or even surpass those of state-of-the-art superconducting SPDs.
+The distinguished advantage of the ESPD paradigm is its foundation on non-superconducting components entirely, obviating the demanding cryogenic operational and infrastructural requirements of SNSPDs, and thereby enabling high-performance SPDs to be utilized beyond specialized low-temperature laboratories.
+Our numerical analysis shows that a legacy non-superconducting SPD (with $\text{DE}\approx 59\%$ and $\text{DCR}\approx 10^{-2}$) can be enhanced to achieve $\text{DE}>93\%$ and $\text{DCR}<10^{-9}$, reducing the minimal tolerable channel transmission rate in quantum communication by several orders of magnitude.
+Although physically realizing such a large-scale architecture requires substantial integration and engineering efforts, the ESPD framework represents a paradigm shift, trading resource overhead for operation-environmental simplicity.
+It opens the door to high-performance quantum-optical applications without cryogenic cooling, providing a clear and viable path toward the widespread implementation of quantum optical tasks in condition-constrained environments.
 Declarations
-The author declares no financial or non-financial competing interests.
+The author declares no competing interests.
+Funding Statements
+There is no funding for this work.
 Author Contributions

### v3

@@
-Enhanced Single-Photon Detector: A framework for Superconducting-Level Performance without cryogenic cooling
+A Theoretical Framework for Iteratively Enhanced Room-Temperature Single-Photon Detection
 Abstract
-High-performance single-photon detectors (SPDs) are indispensable components for a wide range of quantum optical applications.
-However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and necessitates challenging operational conditions, such as cryogenics, thereby hindering scalable implementation.
-To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of conventional room-temperature SPDs.
-Relying entirely on non-superconducting components and eliminating cryogenic cooling requirements, the ESPD scheme can upgrade a legacy SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) of $10^{-2}$, to a device with superior performance metrics, achieving DE exceeding $93\%$ and DCR below $10^{-9}$.
-This performance rivals or even surpasses that of state-of-the-art superconducting SPDs, allowing the minimal tolerable channel transmission rate for quantum key distribution (QKD) protocols to be reduced by several orders of magnitude.
-Although the architecture requires substantial integration effort, the scheme can provide superconducting-level performance without cryogenic cooling, offering a clear path toward the widespread deployment of high-performance SPDs as well as related quantum technologies in infrastructure-constrained environments.
-keywords:
-Room-Temperature Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
-1 Introduction
-Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurements are indispensable for information readout across virtually all applications.
+High-performance photon detection is indispensable in a wide range of quantum-optical applications and is conventionally treated as a fixed device-level operation based on single-photon detectors (SPDs).
+However, state-of-the-art SPDs rely on superconducting materials, which impose severe technological demands and require challenging operational conditions such as cryogenic cooling, thereby hindering scalable implementation.
+Here, we propose the enhanced single-photon detector (ESPD) framework, a theoretical paradigm supported by numerical simulations, that reformulates single-photon detection as an iteratively enhanced process based on state preparation, controlled operations, projective measurements, and multi-copy analysis, and enables substantial performance improvement using exclusively room-temperature components.
+Numerical simulations indicate that, within a physically motivated parameter regime, the ESPD framework can upgrade a legacy SPD with a detection efficiency (DE) of about 59% and a dark count rate (DCR) of about $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$, comparable to those of superconducting SPDs.
+As a consequence, the minimal tolerable channel transmission rate for quantum key distribution protocols can be reduced by several orders of magnitude.
+While its physical realization would require substantial experimental integration, the ESPD framework establishes a general system-level perspective on photon detection, highlighting the potential of iterative quantum processing for overcoming intrinsic detector limitations at room temperature.
+preprint: APS/123-QED
+I Introduction
+Single-photon detection plays a central role in quantum optics and underlies a wide range of quantum-optical applications.
+Conventionally, the detection process is treated as a fixed device-level operation implemented by single-photon detectors (SPDs).
 The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$.
@@
 A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
-Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
-Year Researchers (et al.) Operation temperature Detection efficiency Dark count rate
+Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs. OT: Operation Temperature, DE: Detective efficiency, DCR: Dark count rate.
+Year Researchers (et al.) OT DE DCR
 2020 P. Hu   [30] 2.10 K 95.0 % $0.5\times 10^{-5}$
@@
 This requirement introduces significant infrastructural requirements and operational complexity, fundamentally hindering the scalability and widespread implementation of high-performance quantum technology in field-deployable and environmentally constrained situations.
-To overcome these limitations, we propose the enhanced SPD (ESPD) architecture that can meet advanced performance requirements, potentially outperforming superconducting ones, while relying solely on non-superconducting components suitable for room-temperature operation.
-The proposed method iteratively upgrades the performance of an existing conventional SPD level by level, ultimately to a high-performance one.
-As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] with DE about 59% and DCR of $10^{-2}$ to one achieving DE exceeding 93% and DCR below $10^{-9}$, thereby can rival the state-of-the-art devices and reduce the minimal tolerable channel transmission rate in QKD by several orders of magnitude.
-Despite the substantial engineering integration effort required for its large-scale architecture, the ESPD method offers two key advancements: It not only provides a high-performance SPD design approach without cryogenic components, enabling a practical route toward accurate room-temperature photonic detection, but also establishes a paradigm shift for quantum information processing by leveraging multi-copy analysis and iterative enhancement to boost performance using conventional components.
-Consequently, this work opens a new avenue for the scalable deployment of high-performance SPDs and advances broader quantum optical technologies in infrastructure-constrained environments.
-2 Result
+To address the limitations imposed by the suboptimal performance of room-temperature SPDs and cryogenic operation in superconducting alternatives, we propose the enhanced single-photon detector (ESPD) framework, a system-level theoretical paradigm supported by numerical simulations for improving single-photon detection using exclusively room-temperature components.
+The ESPD framework reformulates single-photon detection as an iteratively enhanced process rather than a fixed device-level operation, in which detection performance is progressively improved through the integration of state preparation, controlled operations, projective measurements, and multi-copy analysis.
+As a representative numerical benchmark, we show that, within a physically motivated parameter regime, the ESPD framework can enhance a legacy non-superconducting SPD reported two decades ago [28], upgrading an initial detector with DE of about 59% and DCR of $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$.
+These values are comparable to those of state-of-the-art superconducting SPDs and relax the minimal tolerable channel transmission rate for quantum key distribution by several orders of magnitude.
+While the physical realization of such an architecture would require substantial experimental integration, the ESPD framework provides a general platform for the theoretical and numerical investigation of enhancement mechanisms in room-temperature photon detection, highlighting the way in which system-level quantum processing can overcome intrinsic device limitations.
+II Method and Result
 In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance initialized from a base (existing) SPD.
 We establish the theoretical framework governing its operation, provide analytical approximations to characterize its performance, and state its applications in QKD.
-2.1 Design of the Enhanced Single-Photon Detector
+II.1 Paradigm of the Enhanced Single-Photon Detection
 The ESPD architecture is designed to recursively upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
@@
 The iteration begins with an initial existing SPD, $\text{ESPD}_{0}$, characterized by initial value $\eta_{0}=\eta$ and $d_{0}=d$.
-The design utilizes qubit logic. The schematic of the ESPD paradigm is illustrated in Fig. 1.
+With qubit logic, the schematic of the ESPD paradigm is illustrated in Fig. 1.
 Figure 1: Schematic representation of the recursive ESPD architecture.
@@
 Therefore, the decision logic is essentially that the positive report rate (among all effective reports) on the multiple detections reaches a pre-defined threshold..
-2.2 Theoretical Formulations
-2.2.1 Notations
+Physical intuition: Intuitively, each level of the ESPD framework increases the number of effective detection opportunities for an input signal, thereby enhancing the overall detection capability.
+In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level generates $n_{s+1}$ identical copies of the incoming signal by the controlled modules, which are independently detected by $n_{s+1}$ SPDs.
+A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered.
+Under this idealized assumption, the detection efficiency can be enhanced by choosing a small threshold $k_{s+1}$, for example, when $k_{s+1}=1$, the effective DE becomes $1-(1-\eta_{s})^{n_{s+1}}$.
+Furthermore, the DCR is strongly suppressed, since a false positive outcome occurs only when dark-count events arise in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$, which decreases dramatically when $k_{s+1}$ is increased.
+This physical intuition remains valid beyond the idealized limit.
+When imperfections in the auxiliary components introduce errors that are sufficiently small relative to the activation probability of a genuine input signal, the ESPD framework continues to improve detection performance.
+A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section II.2.
+II.2 Theoretical Formulations
+The rigorous theoretical formulation of the ESPD framework is developed in this subsection. We first introduce the notations and assumptions, and then present the exact analytical expressions.
+II.2.1 Notations and Basic Assumptions
 We define the parameters governing the system dynamics.
@@
 The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
-For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$.
+For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$, and they are implemented in a feedforward controlled configuration, which ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$.
+However, we emphasize that the generalized ESPD paradigm remains valid for post-selection controlled gates (for a brief discussion, see Section III.1.2).
 Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$).
@@
 The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
-2.2.2 Assumptions
+II.2.2 Parameter Approximation Assumptions
+The assumptions below are employed solely to derive analytical approximations in Section II.3, as well as the approximation part in Eq.
+(1) and (2).
+They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work.
+Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section III.1.
 In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$.
@@
 We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant.
-Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section 3.1).
-Moreover, for analytical simplicity, we assume the controlled gates are implemented in a feedforward controlled configuration.
-This ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$.
-However, we emphasize that the generalized ESPD design remains valid for post-selection controlled gates (for a brief discussion, see Section 3.1.2).
+Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section III.1).
 Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
-2.2.3 Formulations
+II.2.3 Formulations
 Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector.
 They can be expressed as:
-$P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
-$Q_{s+1}=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$ (2)
+$P_{s+1}$ $=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}$ (1)
+$=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$
+$Q_{s+1}$ $=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}$ (2)
+$=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$
 Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final controlled gate, respectively, then:
@@
 If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
-$P_{s+1,k_{s+1},i}$ (4)
-$=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
+$P_{s+1,k_{s+1},i}=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$ (4)
 $+Q^{\prime}_{s+1}\sum_{j_{1}+j_{2}=k_{s+1}-1}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
@@
 If the photon survives all $C(F)$ operations, the probability of a final positive report is:
-$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
-$\qquad\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
-$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
-$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
+$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
+$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
 Combining these, the DE of $\text{ESPD}_{s+1}$ is:
@@
 $d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
-$\qquad\qquad\qquad\qquad\qquad\qquad\quad+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
+$+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
 Collectively, the DE and DCR evolution is governed by the non-linear time-variant dynamic system:
@@
 where the map $G_{s+1}$ is defined by Eq. (1) to (7).
-2.3 Analytical Approximations and Stability Analysis
+II.3 Analytical Approximations and Stability Analysis
 While Eq.
 (8) describes the exact dynamics, an analytical solution is generally intractable.
-However, employing the approximations in Eq.(1) to (3) and the assumptions from Section 2.2.2, we can derive simplified bounds to elucidate the system’s behavior.
+However, employing the approximations in Eq.(1) to (3) and the assumptions from Section II.2.2, we can derive simplified bounds to elucidate the system’s behavior.
 Consider the DCR evolution. Eq. (7) can be reformulated as:
 $d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}$ (9)
-$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
+$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
 Proposition 1.
@@
 $\frac{d}{dx}f(x)$ (11)
-$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$
-$\qquad\qquad\qquad\quad+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
+$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
 $=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}(k-1-nx)+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}(j-nx)]$
@@
 Applying the upper bound from Eq. (2), substituting Eq. (3) to Eq. (9), and assuming $k-1\geq nx$, we obtain:
-$d_{s+1}\leq d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}$ (12)
-$\qquad\qquad\qquad\qquad\qquad\qquad\quad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$
+$d_{s+1}\leq$ $d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$ (12)
 which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate $n_{s+1},k_{s+1}$, we assume $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
-$d_{s+1}$ $\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}$ (13)
-$=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$
+$d_{s+1}\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$ (13)
 This provides an estimation for the DCR at level $s+1$.
 Similarly, the DE can be lower-bounded by:
-$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (14)
-$\qquad\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
-$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
-$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
-$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}$
-$\qquad\qquad\qquad\qquad\qquad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
+$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$ (14)
+$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
+$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
 which is independent of $d_{s}$. Therefore, solving the inequality
@@
 (15).
-2.4 Performance Simulations
-While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
-In this section, we present numerical simulations to validate the ESPD paradigm, demonstrating that substantial performance gains are achievable even with moderate initial hardware specifications.
-For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
-Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs.
-The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
-Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation (see Section 3.1).
-We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section 3.1).
-Performance Outcomes: Achievable performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig.
-2 and 3.
-The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs.
-Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 2: Performance for initial parameters $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
-Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
+II.4 Performance Simulations
+Table 2: Performance of an ESPD initialized with a SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
 Para 1 Para 2 Para 3
@@
 8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 3: Performance for initial parameters $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
-Table 3: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
+Table 3: Performance of an ESPD initialized with a SPD from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
 Para 1 Para 2 Para 3
@@
 8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-2.5 Application on QKD
+While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
+In this section, we present numerical simulations for the ESPD paradigm, indicating that substantial performance gains are achievable even with moderate initial hardware specifications.
+For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
+Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs.
+The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
+Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section III.1 for parameter discussions).
+We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section III.1).
+Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig.
+2 and 3.
+The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs.
+Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 2: Performance for initial parameters $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 3: Performance for initial parameters $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
+II.5 Application on QKD
 To illustrate the practical impact of the ESPD scheme, we consider its application in QKD.
@@
 Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
-Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency.
-This allows significant QKD performance improvements without additional efficiency costs.
-Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
-3 Discussion
-We now discuss the technological feasibility, as well as limitations and avenues for future work.
-3.1 Technological Feasibility and Performance under Degraded Parameters
-The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs.
-The device requirements are discussed as follows.
-3.1.1 State Preparations and Measurements
-The flexibility of the architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one.
-In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates.
-Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
-3.1.2 Controlled Gates
-The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
-A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
-It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module, and consequently, does not decrease with an increasing number of cascaded modules or cascaded levels.
-The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
-However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations.
-While the ESPD design remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded.
-Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$.
-Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate.
-Table 4 and Fig.4 demonstrate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.).
-This result indicates robustness, validating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
-(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
-(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
-Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
-Table 4: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.80 to 0.96.
-Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section 2.4.
+Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96.
+Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section II.4.
 pp=0.80 pp=0.84 pp=0.88 pp=0.92 pp=0.96
@@
 8 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
-We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate.
-Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths.
-Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38].
-Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42].
-However, it is also worth testing the effectiveness of the design under degraded parameters to verify its suitability for standard fabrication processes with lower technology requirements.
-Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6).
-All other settings remain consistent with those described in Section 2.4.
-The numerical results confirm the robustness of the scheme: Comparison of Table 5 to Table 2 demonstrates that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
-This demonstrated robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
+Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency.
+This allows significant QKD performance improvements without additional efficiency costs.
+Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
+III Discussion
+In the absence of experimental implementation, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations and directions for future work.
+III.1 Technological Feasibility and Performance under Degraded Parameters
+The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs.
+The device requirements are discussed as follows.
+III.1.1 State Preparations and Measurements
+The flexibility of the architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one.
+In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates.
+Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
+III.1.2 Controlled Gates
+The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
+(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
+(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
+Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section II.4.
 Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
@@
 8 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 6: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
 Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under a low controlled gate fidelity, $P=0.40$.
@@
 8 (6,2) $(75.4\%,\ 3.7\times 10^{-5})$ (7,2) $(83.9\%,\ 6.4\times 10^{-5})$ (8,2) $(87.7\%,\ 9.6\times 10^{-5})$
-3.1.3 Choice of initialized SPDs
+A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
+It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module, and consequently, does not decrease with an increasing number of cascaded modules or cascaded levels.
+The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
+However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations.
+While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded.
+Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$.
+Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section II.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate.
+Table 4 and Fig.4 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.).
+This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
+Figure 6: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
+We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate.
+Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths.
+Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38].
+Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42].
+However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements.
+Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6).
+All other settings remain consistent with those described in Section II.4.
+The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
+This robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
+III.1.3 Choice of initialized SPDs
 It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation.
-As demonstrated in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval.
+As supported in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval.
 The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point.
@@
 Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient.
-Indeed, simulations confirm that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
+Indeed, simulations suggest that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
 In summary, the ESPD scheme only requires a normal, room-temperature SPD for initialization, which is routinely achievable. For instance, even an SPD developed two decades ago [28] is sufficient.
-3.1.4 Auxiliary Sources
+III.1.4 Auxiliary Sources
 Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, we note that implementing the ESPD scheme may not require single photon sources.
@@
 However, the formulation of this, including its impact on controlled operations, needs a rigorous discussion and is reserved for future work.
-3.2 Limitations and Future Works
-Despite the significance of the ESPD framework, some limitations remain, which are left for future investigations.
-3.2.1 Theoretical Completeness
-The theoretical framework established in Section 2.2 remains an area for deeper exploration.
-While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest.
+III.2 Limitations and Future Works
+Despite the significance of the ESPD framework, some limitations remain and warrant future investigation.
+III.2.1 Theoretical Completeness
+The theoretical framework established in Section II.2 remains an area for deeper exploration.
+While the numerical designs presented in Section II.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest.
 Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required.
 Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
-3.2.2 Resource Overhead vs. Environmental Simplicity
-While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels.
-Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections.
-Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD.
-For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components.
-While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
-However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits.
-Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57].
-Notably, very-large-scale integration of quantum photonics has been experimentally realized with over 2,500 components on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes.
-Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms.
-For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
-In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation.
-The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
-The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration.
-By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost.
-This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
-3.2.3 Optimization Design
-The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments.
-Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$).
-Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence.
-As shown in Section 2.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
-Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Fig.2.
-Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
+III.2.2 Resource Overhead vs. Environmental Simplicity
+Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section II.4
 Para 4 Para 5 Para 6
@@
 8 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
+While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels.
+Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections.
+Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD.
+For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components.
+While not posing a fundamental physical barrier (as discussed in Section III.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
+However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits.
+Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57].
+Notably, very-large-scale integration of quantum photonics has been experimentally realized with over 2,500 components on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes.
+Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms.
+For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
+In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation.
+The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
+The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration.
+By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost.
+This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
+III.2.3 Optimization Design
+The inherent exponential resource scaling discussed in Section III.2.2 necessitates optimization for future practical deployments.
+Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$).
+Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence.
+As shown in Section II.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
+Furthermore, as shown in Table 7 and visualized in Fig.7, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
+Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Fig.2.
 The optimization problem also encompasses the optimal circuit topology.
 To mitigate cumulative error propagation, minimizing the maximal circuit length is essential.
-As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the design: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1.
+As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section II.1.
 However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
-3.2.4 Physical Integrations and Experiments
-Despite the convincing demonstration of effectiveness via numerical simulations in Section 2.4, physical implementations of the ESPD scheme require dedicated experimental validations.
-Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization.
+III.2.4 Physical Integrations and Experiments
+Despite the convincing indication of effectiveness via numerical simulations in Section II.4, physical implementations of the ESPD scheme require dedicated experimental validations.
+Reassuringly, as established in Section III.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization.
 However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components.
@@
 Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
-3.2.5 Count Rate and Timing Jitter
+III.2.5 Count Rate and Timing Jitter
 While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors.
@@
 Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
-4 Conclusion
-We have introduced the Enhanced Single-Photon Detector (ESPD) framework, a robust methodology for upgrading low-performance SPDs into devices whose capabilities can match or even surpass those of state-of-the-art superconducting SPDs.
-The distinguished advantage of the ESPD paradigm is its foundation on non-superconducting components entirely, obviating the demanding cryogenic operational and infrastructural requirements of SNSPDs, and thereby enabling high-performance SPDs to be utilized beyond specialized low-temperature laboratories.
-Our numerical analysis shows that a legacy non-superconducting SPD (with $\text{DE}\approx 59\%$ and $\text{DCR}\approx 10^{-2}$) can be enhanced to achieve $\text{DE}>93\%$ and $\text{DCR}<10^{-9}$, reducing the minimal tolerable channel transmission rate in quantum communication by several orders of magnitude.
-Although physically realizing such a large-scale architecture requires substantial integration and engineering efforts, the ESPD framework represents a paradigm shift, trading resource overhead for operation-environmental simplicity.
-It opens the door to high-performance quantum-optical applications without cryogenic cooling, providing a clear and viable path toward the widespread implementation of quantum optical tasks in condition-constrained environments.
+IV Conclusion
+We have introduced the enhanced single-photon detector (ESPD) framework, which reformulates single-photon detection as an iteratively enhanced process based on integrating state preparations, controlled operations, projective measurements, and multi-copy analysis, rather than as a fixed device-level operation.
+Within this system-level theoretical framework, detector performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
+Numerical simulations indicate that, starting from a conventional non-superconducting SPD, the ESPD framework can substantially enhance DE while strongly suppressing DCR, yielding effective performance metrics comparable to those of state-of-the-art superconducting SPDs, and significantly relaxing the minimal tolerable channel transmission rate in quantum communication.
+The ESPD framework makes explicit the trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure with circuit integration complexity.
+While the experimental realization of large-scale ESPD architectures will require substantial engineering effort, the parameter studies presented in this work, grounded in the performance of existing photonic components, support the in-principle viability of the framework.
+In essence, this work establishes a system-level perspective on photon detection, showing that intrinsic device limitations can be transcended through iterative quantum processing and providing a principled foundation for the development of high-performance, room-temperature SPDs.
+Data and Code Availability
+All data/codes generated or used during this study are included in this published article or can be found at https://github.com/Hao-B-Shu/ESPD.
+Author Contributions
+Hao Shu (the unique author): study design, investigation, experiment, analysis, manuscript preparation.
 Declarations
 The author declares no competing interests.
-Funding Statements
-There is no funding for this work.
-Author Contributions
-Hao Shu (the unique author): study design, investigation, experiment, analysis, manuscript preparation.
-Data and Code Availability
-All data/codes generated or used during this study are included in this published article or can be found at https://github.com/Hao-B-Shu/ESPD.

### v4

@@
-A Theoretical Framework for Iteratively Enhanced Room-Temperature Single-Photon Detection
+Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature
 Abstract
-High-performance photon detection is indispensable in a wide range of quantum-optical applications and is conventionally treated as a fixed device-level operation based on single-photon detectors (SPDs).
-However, state-of-the-art SPDs rely on superconducting materials, which impose severe technological demands and require challenging operational conditions such as cryogenic cooling, thereby hindering scalable implementation.
-Here, we propose the enhanced single-photon detector (ESPD) framework, a theoretical paradigm supported by numerical simulations, that reformulates single-photon detection as an iteratively enhanced process based on state preparation, controlled operations, projective measurements, and multi-copy analysis, and enables substantial performance improvement using exclusively room-temperature components.
-Numerical simulations indicate that, within a physically motivated parameter regime, the ESPD framework can upgrade a legacy SPD with a detection efficiency (DE) of about 59% and a dark count rate (DCR) of about $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$, comparable to those of superconducting SPDs.
-As a consequence, the minimal tolerable channel transmission rate for quantum key distribution protocols can be reduced by several orders of magnitude.
-While its physical realization would require substantial experimental integration, the ESPD framework establishes a general system-level perspective on photon detection, highlighting the potential of iterative quantum processing for overcoming intrinsic detector limitations at room temperature.
-preprint: APS/123-QED
-I Introduction
-Single-photon detection plays a central role in quantum optics and underlies a wide range of quantum-optical applications.
-Conventionally, the detection process is treated as a fixed device-level operation implemented by single-photon detectors (SPDs).
-The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$.
-The former affects the efficiency of nearly all quantum operations, while the latter imposes the ultimate limitation in high-attenuation scenarios.
-In terms of DE, numerous quantum experiments rely on SPD signals to retrieve results.
-In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3].
-In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12].
-In heralded single-photon sources, where entangled pairs are generated, and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16].
-With respect to DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD).
-The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, significantly lower than 50% [17, 18, 19, 20].
-However, communication channels induce exponential photon attenuation [21, 22, 23, 24].
-If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics.
-Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
-Driven by their critical importance, SPD has remained a continuous research focus since the inception of quantum technologies.
-While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR.
-For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with DE about 59% and DCR of $10^{-2}$ [28], while an InGaAs/InP SPD presented in 2017 operates at 223 K with DE about 27.5% and DCR of $10^{-6}$ [29].
-Such specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$.
-To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33].
-These SPDs achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude.
-This capability enables the implementation of demanding quantum tasks, such as long-distance QKD [19, 34, 35, 36].
-A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
-Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs. OT: Operation Temperature, DE: Detective efficiency, DCR: Dark count rate.
-Year Researchers (et al.) OT DE DCR
+Photon detection, a fundamental quantum technology, is traditionally treated as a fixed device-level operation governed by intrinsic properties of single-photon detectors (SPDs).
+High-performance detection has therefore largely relied on superconducting technologies whose requirement for cryogenic operation imposes substantial infrastructure constraints, limiting scalable deployments.
+Here, the enhanced single-photon detection (ESPD) framework is presented as a system-level theoretical paradigm that shifts photon detection from device-centric optimization to an integrated quantum-information-processing task, by reformulating it as an iteratively enhanced process integrating state-preparation, controlled operations, projective measurements, and multi-copy analysis.
+ESPD enables systematic performance enhancement through architectural design rather than material modification, thereby circumventing superconducting components, allowing high-performance detection using exclusively room-temperature hardware.
+Numerical simulations grounded in physically motivated parameters indicate that the ESPD framework can upgrade a conventional room-temperature SPD to effective DE exceeding 93% and DCR below $10^{-9}$, which are comparable to state-of-the-art superconducting SPDs and can significantly relax the minimal tolerable channel transmission rate in quantum communications.
+While physical realization would require further component integration, this work establishes a rigorous theoretical framework for enhancing detection performance through architectural quantum-information principles, providing a general blueprint for transcending device-level constraints and guiding the development of next-generation room-temperature quantum technologies.
+keywords:
+Single-photon detection, System-level quantum architecture, Iterative quantum information processing, Room-temperature quantum technologies, Photonic integration
+Hao Shu*
+[…]
+{affiliations}
+Dr. Hao Shu
+Address: Sun Yat-Sen University, Guangzhou, China
+Email Address: Hao_B_Shu@163.com (Primary), shuh6@mail.sysu.edu.cn (Institutional)
+1 Introduction
+Single-photon detection plays a central role in quantum optics and serves as the bedrock for a wide range of quantum-information applications.
+Conventionally, the detection process is treated as a fixed device-level operation primarily determined by the material properties of single-photon detectors (SPDs).
+The primary performance metrics characterizing an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$.
+The former directly dictates the success probability of quantum operations conditioned on detection events, while the latter imposes a fundamental signal-to-noise constraint in high-attenuation scenarios.
+In terms of DE, high-performance detection is indispensable for retrieving meaningful outcomes in quantum experiments.
+In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [1, 2, 3].
+In fundamental tests of quantum nonlocality, verifying violations of the Clauser-Horne-Shimony-Holt or Eberhard inequalities requires DE exceeding strict thresholds to close detection loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12].
+In heralded single-photon sources, where one photon of an entangled pair signals the presence of the other, the achievable heralding rate is also strictly limited by DE [13, 14, 15, 16].
+With respect to DCR, a particularly stringent constraint arises in quantum communication, notably in quantum key distribution (QKD).
+The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are significantly lower than 50% [17, 18, 19, 20].
+However, optical communication channels inevitably induce exponential photon attenuation with distance [21, 22, 23, 24].
+When the effective signal rate, determined by channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts, which contribute a 50% error rate, causing the QBER to rise precipitously.
+Consequently, the achievable secure distance of QKD systems is ultimately bounded by the DCR of the employed SPDs [25, 26, 27].
+Motivated by these critical demands, the development of high-performance SPDs has remained a primary research focus.
+While conventional semiconductor-based SPDs offer the advantage of non-cryonic operation, they typically suffer from an intrinsic trade-off between limited DE and relatively high DCR.
+For instance, a frequency upconversion SPD demonstrated in 2004 operated at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operated at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29].
+Such specifications often fall short of the requirements for advanced quantum protocols, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
+To meet these stringent requirements, substantial effort has shifted toward designing SPDs with superconducting technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33].
+These devices have successfully pushed DE beyond 90% while suppressing DCR by several orders of magnitude compared to semiconductor counterparts, thereby enabling demanding applications such as long-distance QKD [19, 34, 35, 36].
+Recent representative benchmarks of state-of-the-art superconducting SPDs are summarized in Table 1.
+Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
+Year Researchers (et al.) Operation Temperature Detection Efficiency Dark Count Rate
 2020 P. Hu   [30] 2.10 K 95.0 % $0.5\times 10^{-5}$
@@
 2024 I. Charaev [37] 20.0 K 7.60 % $1.0\times 10^{-3}$
-However, as illustrated in the table, the reliance on superconductors necessitates stringent operating conditions, specifically with extremely low temperatures.
-When these demanding conditions are relaxed, the SPD performance degrades significantly.
-This requirement introduces significant infrastructural requirements and operational complexity, fundamentally hindering the scalability and widespread implementation of high-performance quantum technology in field-deployable and environmentally constrained situations.
-To address the limitations imposed by the suboptimal performance of room-temperature SPDs and cryogenic operation in superconducting alternatives, we propose the enhanced single-photon detector (ESPD) framework, a system-level theoretical paradigm supported by numerical simulations for improving single-photon detection using exclusively room-temperature components.
-The ESPD framework reformulates single-photon detection as an iteratively enhanced process rather than a fixed device-level operation, in which detection performance is progressively improved through the integration of state preparation, controlled operations, projective measurements, and multi-copy analysis.
-As a representative numerical benchmark, we show that, within a physically motivated parameter regime, the ESPD framework can enhance a legacy non-superconducting SPD reported two decades ago [28], upgrading an initial detector with DE of about 59% and DCR of $10^{-2}$ to effective performance metrics with DE exceeding 93% and DCR below $10^{-9}$.
-These values are comparable to those of state-of-the-art superconducting SPDs and relax the minimal tolerable channel transmission rate for quantum key distribution by several orders of magnitude.
-While the physical realization of such an architecture would require substantial experimental integration, the ESPD framework provides a general platform for the theoretical and numerical investigation of enhancement mechanisms in room-temperature photon detection, highlighting the way in which system-level quantum processing can overcome intrinsic device limitations.
-II Method and Result
-In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance initialized from a base (existing) SPD.
-We establish the theoretical framework governing its operation, provide analytical approximations to characterize its performance, and state its applications in QKD.
-II.1 Paradigm of the Enhanced Single-Photon Detection
-The ESPD architecture is designed to recursively upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
-For clarity, we clarify that DE is the probability of generating a positive detection event given a non-vacuum (single-photon) input, and the DCR is the probability of a positive report given a vacuum input.
-Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with DE $\eta_{s}$ and DCR $d_{s}$.
-The iteration begins with an initial existing SPD, $\text{ESPD}_{0}$, characterized by initial value $\eta_{0}=\eta$ and $d_{0}=d$.
+However, the superior performance of superconducting SPDs comes at the cost of cryogenic operation.
+This imposes stringent environmental requirements and necessitates bulky, power-intensive cooling infrastructure.
+The significant degradation of performance at elevated temperatures highlights a strong trade-off between detection performance and operating conditions.
+These constraints present major barriers to scalable deployment, particularly in field applications, satellite payloads, and scenarios with strict size, weight, and power limitations where cryogenic infrastructure is impractical.
+To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, the enhanced single-photon detection (ESPD) framework is introduced, shifting the focus from device-centric optimization to a system-level architectural paradigm.
+Within this framework, single-photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on the material properties of a single SPD.
+Through structured system-level integration, the ESPD framework enables effective enhancement of detection performance using exclusively room-temperature components.
+Numerical simulations, grounded in physically motivated parameters drawn from commercially available devices and experimentally reported SPDs, indicate that, under an exclusively room-temperature component setting, the ESPD framework enables effective DE exceeding 93% together with effective DCR below $10^{-9}$, starting from a legacy SPD (for example, the one reported two decades ago with about 59% DE and $10^{-2}$ DCR [28]).
+These effective performance metrics are comparable to those of state-of-the-art superconducting SPDs and lead to a substantial relaxation of the minimal tolerable channel transmission rate in QKD.
+While physical implementation would require further system-level integration of quantum technologies, the ESPD framework provides a structured theoretical basis for achieving high-performance photon detection at room temperature.
+By explicitly analyzing technological feasibility and the trade-off between resource overhead and operating temperature, the ESPD framework offers a feasible technological roadmap and practical guidance for future experimental efforts.
+It serves as a theoretical blueprint for overcoming intrinsic device limitations, with the potential to unlock a wide range of quantum applications in environmentally constrained scenarios.
+2 Method and Result
+This section details the architectural structure of the ESPD framework and formulates the corresponding detection process.
+The ESPD paradigm is introduced in Section 2.1, with theoretical formulations established in Section 2.2.
+Analytical approximations are developed in Section 2.3, and numerical simulations are presented in Section 2.4.
+Finally, implications for QKD are discussed in Section 2.5.
+2.1 Paradigm of the Enhanced Single-Photon Detection
+The ESPD architecture is formulated as a level-by-level framework in which the effective detection performance of a given SPD is enhanced through successive processing stages.
+Here, the enhancement is quantified in terms of DE and DCR, defined respectively as the probability of generating a positive detection event given a non-vacuum (single-photon) input and the probability of a positive report given a vacuum input.
+Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with effective DE $\eta_{s}$ and effective DCR $d_{s}$.
+The iteration begins with an existing SPD, denoted as $\text{ESPD}_{0}$, characterized by the initial values $\eta_{0}=\eta$ and $d_{0}=d$.
 With qubit logic, the schematic of the ESPD paradigm is illustrated in Fig. 1.
@@
 Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective measurements via basis $\{|0\rangle,|1\rangle\}$, and filters represent that the $|0\rangle$ path is aborted.
-The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, proceeds as follows.
-To facilitate understanding, readers can consider an implementation where the degree of freedom (DOF) is photon polarization and the controlled operations are C-NOT gates.
+The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, is defined as follows.
+As a representative example, an implementation may be considered in which the degree of freedom (DOF) is photon polarization and the controlled operations correspond to controlled-NOT (C-NOT) gates.
   1. 1.
-    DOF Selection: Select a manipulable DOF of the incoming photon (e.g., polarization).
-Let $\{|0\rangle,|1\rangle\}$ denote the computational basis for this DOF.
-We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon [^1] Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum, which degenerates to the qubit case.
-However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from photon existence..
+    DOF Selection: A manipulable DOF of the incoming photon is chosen (for example, polarization).
+The computational basis for this DOF is denoted by $\{|0\rangle,|1\rangle\}$.
+The signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit.
+In practical architectures, an internal DOF such as polarization, distinct from photon existence, is often more convenient..
   2. 2.
-    State Preparation: Apply a state preparation operation to the incoming signal to initialize the selected DOF to $|1\rangle$.
-Consequently, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum signal remains $|\emptyset\rangle$.
+    State Preparation: A state-preparation operation is applied to initialize the selected DOF to $|1\rangle$.
+As a result, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum input remains in state $|\emptyset\rangle$.
   3. 3.
-    Controlled Operations: Implement $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates).
-These gates interact the signal path with $n_{s+1}$ auxiliary sources, each initialized in the state $|0\rangle$.
+    Controlled Operations: A set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), are applied to couple the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
   4. 4.
-    Auxiliary Measurement: Perform projective measurements on the auxiliary outputs in the basis $\{|0\rangle,|1\rangle\}$.
-Outcome paths of $|0\rangle$ are discarded, and only the $|1\rangle$ paths are preserved for detection.
+    Auxiliary Measurement: Projective measurements are performed on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$.
+Outcome paths corresponding to $|0\rangle$ are discarded, and only the $|1\rangle$ outcomes are retained for subsequent detection.
   5. 5.
-    Detection: Detect the signals at the end of paths, including the signal one and the auxiliary ones, using $s$-th level detectors, $\text{ESPD}_{s}$.
+    Detection: Detection is carried out on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
   6. 6.
-    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detections are registered among the $n_{s+1}+1$ detectors (comprising $n_{s+1}$ auxiliary detectors and one signal path detector).
+    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path.
 Otherwise, a negative report is returned.
-Here, $k_{s+1}$ is a pre-selected threshold integer, and $n_{s+1}$ refers to the number of effective operations [^2] For example, if the controlled gates are deterministic, then $n_{k+1}$ is exactly the number of used operations, while if the controlled gates are probabilistic, then $n_{k+1}$ refers to the successful ones after post-selection.
-Therefore, the decision logic is essentially that the positive report rate (among all effective reports) on the multiple detections reaches a pre-defined threshold..
-Physical intuition: Intuitively, each level of the ESPD framework increases the number of effective detection opportunities for an input signal, thereby enhancing the overall detection capability.
-In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level generates $n_{s+1}$ identical copies of the incoming signal by the controlled modules, which are independently detected by $n_{s+1}$ SPDs.
+Here, $k_{s+1}$ is a preselected threshold integer, and $n_{s+1}$ denotes the number of effective controlled operations [^2] The main discussions of the article assume deterministic controlled gates, for which $n_{s+1}$ equals the number of implemented operations.
+For probabilistic gates, $n_{s+1}$ refers to the number of successful operations after post-selection.
+In this case, the decision logic is naturally reformulated in terms of the fraction of positive detection outcomes among the effective detections exceeding a predefined threshold..
+Physical intuition: From a detection perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal.
+In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ effective signal instances through the controlled modules, which are independently detected by $n_{s+1}$ SPDs.
 A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered.
-Under this idealized assumption, the detection efficiency can be enhanced by choosing a small threshold $k_{s+1}$, for example, when $k_{s+1}=1$, the effective DE becomes $1-(1-\eta_{s})^{n_{s+1}}$.
-Furthermore, the DCR is strongly suppressed, since a false positive outcome occurs only when dark-count events arise in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$, which decreases dramatically when $k_{s+1}$ is increased.
-This physical intuition remains valid beyond the idealized limit.
-When imperfections in the auxiliary components introduce errors that are sufficiently small relative to the activation probability of a genuine input signal, the ESPD framework continues to improve detection performance.
-A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section II.2.
-II.2 Theoretical Formulations
-The rigorous theoretical formulation of the ESPD framework is developed in this subsection. We first introduce the notations and assumptions, and then present the exact analytical expressions.
-II.2.1 Notations and Basic Assumptions
-We define the parameters governing the system dynamics.
+Under this idealized assumption, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$.
+Furthermore, the effective DCR is suppressed, since a false positive outcome requires dark-count events to occur in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$.
+This physical intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal.
+A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section 2.2.
+2.2 Theoretical Formulations
+The rigorous theoretical formulation of the ESPD framework is developed in this subsection.
+The relevant notations and assumptions are introduced first, followed by the exact analytical expressions and the approximation analysis.
+2.2.1 Notations and Basic Assumptions
+The parameters governing the system behavior are defined as follows.
 As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively.
-Although $\eta_{s}$ may depend on the photon number distribution, we approximate it as a unified value for the given system for simplicity.
+Although $\eta_{s}$ may depend on the photon number distribution, it is approximated as a unified value for the given system for simplicity.
 The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
-For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$, and they are implemented in a feedforward controlled configuration, which ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$.
-However, we emphasize that the generalized ESPD paradigm remains valid for post-selection controlled gates (for a brief discussion, see Section III.1.2).
+For convenience, all controlled operations $C(F_{i})$ are assumed to be identical and are denoted simply as $C(F)$.
+They are implemented in a feedforward controlled configuration, such that $n_{s+1}$ represents both the number of effective modules and the total number of physical modules required at level $s+1$.
+However, the generalized ESPD paradigm remains valid for post-selection controlled gates (see Section 3.1.2 for a brief discussion).
 Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$).
@@
 Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection).
-We approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ module operations.
+The parameters $p$, $P$, and $Q$ are approximately assumed to be invariant across all $C(F)$ module operations.
 It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module.
@@
 The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
-II.2.2 Parameter Approximation Assumptions
-The assumptions below are employed solely to derive analytical approximations in Section II.3, as well as the approximation part in Eq.
-(1) and (2).
+2.2.2 Parameter Approximation Assumptions
+The assumptions below are employed solely to derive analytical approximations in Section 2.3, as well as the approximation part in Eq.
+(1) to (3).
 They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work.
-Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section III.1.
+Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section 3.1.
 In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$.
 Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds.
-We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant.
-Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section III.1).
-Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
-II.2.3 Formulations
-Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector.
+It is further assumed that $d_{s}\ll P\eta_{s}$: If this condition were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant.
+Additionally, the transmission rate of a single controlled module is assumed to satisfy $p\approx 1$, and the SPAM error of the auxiliary path is assumed to satisfy $Q\ll 1$, consistent with the maturity of commercial optical components (see Section 3.1).
+Finally, considering practical implementation costs, $n_{s}$ and $k_{s}$ are taken to be moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
+2.2.3 Formulations
+Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum and vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector.
 They can be expressed as:
-$P_{s+1}$ $=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}$ (1)
-$=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$
-$Q_{s+1}$ $=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}$ (2)
-$=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$
-Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final controlled gate, respectively, then:
+$P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
+$Q_{s+1}$ $=Q[\eta_{s}+(1-\eta_{s})d_{s}]+(1-Q)d_{s}=Q\eta_{s}(1-d_{s})+d_{s}\approx Q\eta_{s}+d_{s}\leq Q+d_{s}$ (2)
+Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum and vacuum output from the final controlled gate, respectively, then:
 $P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\ Q^{\prime}_{s+1}=d_{s}$ (3)
-We now derive the DE and DCR for the $(s+1)$-th level.
-For the DE calculation, we assume a non-vacuum input and calculate the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections.
+The DE and DCR for the $(s+1)$-th level are derived as follows.
+For the DE calculation, a non-vacuum input is assumed, and the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections is evaluated.
 If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
@@
 $=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
-Combining these, the DE of $\text{ESPD}_{s+1}$ is:
+By combining the above contributions, the DE of $\text{ESPD}_{s+1}$ is given by:
 $\eta_{s+1}=p^{n_{s+1}}P_{s+1,k_{s+1}}+\sum_{i=1}^{n_{s+1}}p^{i-1}(1-p)P_{s+1,k_{s+1},i}$ (6)
-For the DCR calculation, we assume a vacuum input. The probability of a false positive is:
-$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
-$+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
+For the DCR calculation, a vacuum input is assumed. The probability of a false positive is given by:
+$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
 Collectively, the DE and DCR evolution is governed by the non-linear time-variant dynamic system:
@@
 where the map $G_{s+1}$ is defined by Eq. (1) to (7).
-II.3 Analytical Approximations and Stability Analysis
+2.3 Analytical Approximations and Stability Analysis
 While Eq.
 (8) describes the exact dynamics, an analytical solution is generally intractable.
-However, employing the approximations in Eq.(1) to (3) and the assumptions from Section II.2.2, we can derive simplified bounds to elucidate the system’s behavior.
-Consider the DCR evolution. Eq. (7) can be reformulated as:
-$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}$ (9)
-$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
+However, by employing the approximations in Eq.
+(1) to (3) together with the assumptions in Section 2.2.2, simplified bounds can be derived to elucidate the system behavior.
+The evolution of the DCR is considered first. Eq. (7) can be reformulated as:
+$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (9)
 Proposition 1.
@@
 $f(x)=a\binom{n}{k-1}x^{k-1}(1-x)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}x^{j}(1-x)^{n-j}$ (10)
-Then $f(x)$ is monotonically increasing for $x\in[0,\frac{k-1}{n}]$.
+for $a\geq 0$, then $f(x)$ is monotonically increasing for $x\in[0,\frac{k-1}{n}]$.
 Proof: For $x\in[0,\frac{k-1}{n}]$:
-$\frac{d}{dx}f(x)$ (11)
-$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
+$\frac{d}{dx}f(x)=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$ (11)
+$+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
 $=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}(k-1-nx)+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}(j-nx)]$
@@
 $\blacksquare$
-Applying the upper bound from Eq. (2), substituting Eq. (3) to Eq. (9), and assuming $k-1\geq nx$, we obtain:
+By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, the following bound is obtained:
 $d_{s+1}\leq$ $d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$ (12)
-which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate $n_{s+1},k_{s+1}$, we assume $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
+which is notably independent of $\eta_{s}$.
+Given $Q_{s},d_{s}\ll 1$ and moderate values of $n_{s+1}$ and $k_{s+1}$, it is assumed that $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$.
+This yields the approximation:
 $d_{s+1}\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$ (13)
@@
 $\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
-which is independent of $d_{s}$. Therefore, solving the inequality
+which is independent of $d_{s}$. Therefore, the condition for DE enhancement at the subsequent level is determined by solving the inequality
 $p^{n}(Px)^{k-1}[x\binom{n}{k-1}(1-Px)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}(Px)^{j-k+1}(1-Px)^{n-j}]-x>0$ (15)
-determines the condition under which the DE is enhanced in the subsequent level.
-If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq.
-(15).
-II.4 Performance Simulations
+If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
+2.4 Numerical Simulations
+While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
+In this section, numerical simulations of the ESPD paradigm are presented, indicating that substantial performance gains can be achieved even with moderate initial hardware specifications.
+For example, a three-level ESPD can upgrade a semiconductor SPD to effective performance metrics with DE $93\%$ and DCR $10^{-9}$, comparable to those of state-of-the-art SNSPDs without reliance on superconducting materials.
+Simulation Parameters: The performance of the ESPD scheme is evaluated using two distinct baseline SPDs.
+The first represents an older one, characterized by parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
+Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section 3.1 for parameter discussions).
+Supplementary simulations with degraded parameter settings are also performed to evaluate robustness in Section 3.1.
+Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig.
+2.
+The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs.
+Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
 Table 2: Performance of an ESPD initialized with a SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
@@
 8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
-In this section, we present numerical simulations for the ESPD paradigm, indicating that substantial performance gains are achievable even with moderate initial hardware specifications.
-For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
-Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs.
-The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
-Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section III.1 for parameter discussions).
-We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section III.1).
-Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig.
-2 and 3.
-The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs.
-Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 2: Performance for initial parameters $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 3: Performance for initial parameters $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
-II.5 Application on QKD
-To illustrate the practical impact of the ESPD scheme, we consider its application in QKD.
-For a QKD protocol having a secure threshold $e_{th}$ and total error rate $e_{C}$ not accounting for SPDs, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
+(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
+(d) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
+Figure 2: Performance of ESPD initialized with SPDs from [28] ((a), (b)) and [29] ((c), (d)), respectively.
+2.5 Implications for QKD
+As a representative application scenario, the implications of the ESPD framework for QKD are analyzed.
+For a QKD protocol characterized by a secure error threshold $e_{th}$ and a total error rate $e_{C}$ excluding detector contributions, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by[27]:
 $\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
-Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude (even compared to SNSPDs), it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor.
-Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
+Since the ESPD paradigm can substantially suppress the effective DCR $d$, the minimal tolerable channel transmission rate $\gamma$ is proportionately reduced.
+Furthermore, the simultaneous enhancement of the DE $\eta$ not only reduces $\gamma$ but also directly improves the achievable key rate at a fixed channel transmission rate.
+Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD framework achieves this improvement without introducing additional sifting procedures that reduce the effective sifting efficiency.
+Here, QKD performance can be enhanced through the simultaneous suppression of the effective DCR and enhancement of the effective DE, without the need for auxiliary sifting operations.
+In this sense, the ESPD paradigm provides a more efficient system-level approach for improving QKD performance under practical constraints.
+3 Discussion
+In the absence of an experimental demonstration, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations as well as directions for future work.
+3.1 Technological Feasibility and Performance under Degraded Parameters
+The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs.
+The device requirements are discussed as follows.
+3.1.1 State Preparations and Measurements
+The flexibility of the ESPD architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one.
+In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be achieved at extremely low levels.
+Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
+3.1.2 Controlled Gates
+The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
 Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96.
-Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section II.4.
+Simulations are performed for $(n,k)=(5,2)$ at each level, and all other settings remain consistent with those described in Section 2.4.
 pp=0.80 pp=0.84 pp=0.88 pp=0.92 pp=0.96
@@
 8 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
-Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency.
-This allows significant QKD performance improvements without additional efficiency costs.
-Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
-III Discussion
-In the absence of experimental implementation, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations and directions for future work.
-III.1 Technological Feasibility and Performance under Degraded Parameters
-The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs.
-The device requirements are discussed as follows.
-III.1.1 State Preparations and Measurements
-The flexibility of the architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one.
-In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates.
-Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
-III.1.2 Controlled Gates
-The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
 (a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
 (b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
-Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section II.4.
+Figure 3: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
+A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
+It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, does not decrease with an increasing number of cascaded modules or ESPD levels.
+The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
+However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations.
+While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded.
+Specifically, if the post-selection efficiency of the controlled gate is $\frac{1}{N}$, the effective average transmission rate from one module to the next is reduced to $p^{N}$.
+Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate.
+Table 4 and Fig.3 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.).
+This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
+We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate.
+Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity should account for the correct operation rate of both the control and controlled paths.
+Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38].
+Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryonic components[39, 40, 41, 42].
+However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements.
+Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), with visual representation in Fig.4.
+All other settings remain consistent with those described in Section 2.4.
+The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
+This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with existing room-temperature quantum technologies.
 Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
@@
 8 (6,2) $(75.4\%,\ 3.7\times 10^{-5})$ (7,2) $(83.9\%,\ 6.4\times 10^{-5})$ (8,2) $(87.7\%,\ 9.6\times 10^{-5})$
-A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
-It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module, and consequently, does not decrease with an increasing number of cascaded modules or cascaded levels.
-The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
-However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations.
-While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded.
-Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$.
-Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section II.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate.
-Table 4 and Fig.4 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.).
-This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 6: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
-We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate.
-Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths.
-Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38].
-Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42].
-However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements.
-Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6).
-All other settings remain consistent with those described in Section II.4.
-The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
-This robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
-III.1.3 Choice of initialized SPDs
-It is also important to note that the requirements for SPD used for initialization are minimal, further lowering the barrier to implementation.
-As supported in comparing Table 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the starting detector, provided it falls within a certain convergence interval.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
+(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.40$.
+(d) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$, with $P=0.40$.
+Figure 4: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$ ((a), (b)) and $P=0.40$ ((c), (d)), respectively.
+3.1.3 Choice of initialized SPDs
+Notably, the requirements imposed on the SPD used for initialization are minimal, which further lowers the practical barrier to implementation.
+As evidenced by the comparison between Tables 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the initialized SPD, provided that it lies within a certain convergence interval.
 The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point.
@@
 Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient.
-Indeed, simulations suggest that both an SPD with low DE and low DCR and one with moderate DE and high DCR serve as viable initialization.
-In summary, the ESPD scheme only requires a normal, room-temperature SPD for initialization, which is routinely achievable. For instance, even an SPD developed two decades ago [28] is sufficient.
-III.1.4 Auxiliary Sources
-Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, we note that implementing the ESPD scheme may not require single photon sources.
+Indeed, numerical simulations suggest that both SPDs with low DE and low DCR[29], as well as those with moderate DE and relatively high DCR[28], can serve as viable initializations.
+In summary, the ESPD scheme requires only a conventional room-temperature SPD for initialization, which is routinely achievable with existing technology.
+For instance, the SPD developed two decades ago [28] is sufficient.
+3.1.4 Auxiliary Sources
+Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single photon sources.
 The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$.
 Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects.
-Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary sources, which will be filtered out and thereby have no effect even if the source is multiple-photon..
-However, the formulation of this, including its impact on controlled operations, needs a rigorous discussion and is reserved for future work.
-III.2 Limitations and Future Works
+Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary paths, which will be filtered out and thereby have no effect even if the source is multiple-photon..
+However, a rigorous formulation of this effect, including its impact on the controlled operations, remains an open problem and is left for future investigation.
+3.2 Limitations and Future Works
 Despite the significance of the ESPD framework, some limitations remain and warrant future investigation.
-III.2.1 Theoretical Completeness
-The theoretical framework established in Section II.2 remains an area for deeper exploration.
-While the numerical designs presented in Section II.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest.
+3.2.1 Theoretical Completeness
+The theoretical framework established in Section 2.2 remains an area for deeper exploration.
+While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest.
 Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required.
-Given the inherent mathematical difficulty of such a non-linear and time-variant system and the sufficiency for practical technology employments by numerical designs, these specialized mathematical studies are reserved for the future.
-III.2.2 Resource Overhead vs. Environmental Simplicity
-Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section II.4
+Given the inherent mathematical difficulty of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
+3.2.2 Resource Overhead vs. Environmental Simplicity
+While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels.
+Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections.
+Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD.
+For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components.
+While not posing a fundamental physical barrier (as discussed in Section 3.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
+However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits.
+Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57].
+Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes.
+Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms.
+For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
+In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation.
+The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
+The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration.
+By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost.
+This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
+3.2.3 Optimization Design
+Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
 Para 4 Para 5 Para 6
@@
 8 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels.
-Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections.
-Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD.
-For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (if not applying techniques such as multiplexing), resulting in hundreds of integrated components.
-While not posing a fundamental physical barrier (as discussed in Section III.1), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
-However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits.
-Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57].
-Notably, very-large-scale integration of quantum photonics has been experimentally realized with over 2,500 components on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes.
-Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms.
-For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
-In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation.
-The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
-The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration.
-By leveraging large-scale photonic platforms, thousands of components can be lithographically fabricated on a single chip, effectively converting high operational costs and environmental constraints into a one-time design and fabrication cost.
-This trade-off is particularly advantageous for applications where size, weight, and power constraints prohibit the use of cryogenics, such as satellite-based quantum communication payloads, mobile quantum nodes, or research laboratories lacking dedicated low-temperature facilities.
-III.2.3 Optimization Design
-The inherent exponential resource scaling discussed in Section III.2.2 necessitates optimization for future practical deployments.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
+Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Table.2.
+The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments.
 Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$).
 Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence.
-As shown in Section II.4, the optimal choice of parameters $n$ (number of auxiliary signals in each level) and $k$ (decision threshold) is crucial: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
+As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
 Furthermore, as shown in Table 7 and visualized in Fig.5, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Fig.2.
 The optimization problem also encompasses the optimal circuit topology.
 To mitigate cumulative error propagation, minimizing the maximal circuit length is essential.
-As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section II.1.
+As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^4] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1.
 However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
-III.2.4 Physical Integrations and Experiments
-Despite the convincing indication of effectiveness via numerical simulations in Section II.4, physical implementations of the ESPD scheme require dedicated experimental validations.
-Reassuringly, as established in Section III.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization.
+3.2.4 Physical Integrations and Experiments
+Despite the convincing evidence provided by the numerical simulations in Section 2.4, the physical implementation of the ESPD scheme still requires dedicated experimental validation.
+Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization.
 However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components.
@@
 Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
-III.2.5 Count Rate and Timing Jitter
+3.2.5 Count Rate and Timing Jitter
 While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors.
 Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices).
-However, the architecture does not impose significant additional penalties: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening.
+However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening.
 Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices.
 Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
-IV Conclusion
-We have introduced the enhanced single-photon detector (ESPD) framework, which reformulates single-photon detection as an iteratively enhanced process based on integrating state preparations, controlled operations, projective measurements, and multi-copy analysis, rather than as a fixed device-level operation.
-Within this system-level theoretical framework, detector performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
-Numerical simulations indicate that, starting from a conventional non-superconducting SPD, the ESPD framework can substantially enhance DE while strongly suppressing DCR, yielding effective performance metrics comparable to those of state-of-the-art superconducting SPDs, and significantly relaxing the minimal tolerable channel transmission rate in quantum communication.
-The ESPD framework makes explicit the trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure with circuit integration complexity.
-While the experimental realization of large-scale ESPD architectures will require substantial engineering effort, the parameter studies presented in this work, grounded in the performance of existing photonic components, support the in-principle viability of the framework.
-In essence, this work establishes a system-level perspective on photon detection, showing that intrinsic device limitations can be transcended through iterative quantum processing and providing a principled foundation for the development of high-performance, room-temperature SPDs.
-Data and Code Availability
+4 Conclusion
+In summary, the enhanced single-photon detection (ESPD) framework is established to reformulate photon detection as an iteratively enhanced quantum-information-processing task composed of state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a fixed device-level operation.
+Within this system-level theoretical framework, detection performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
+Numerical simulations, grounded in experimentally reported parameters and commercially available photonic components, demonstrate that the ESPD framework can substantially enhance the effective detection efficiency (DE) (from 59% to 93%) while strongly suppressing the effective dark count rate (DCR) (from $10^{-2}$ to $10^{-9}$) even when initialized with a conventional room-temperature SPD reported two decades ago.
+The resulting effective performance metrics are comparable to those of state-of-the-art superconducting single-photon detectors, significantly relaxing the minimal tolerable channel transmission rate in quantum communication applications.
+Beyond performance enhancement, the ESPD framework makes explicit the fundamental trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure requirements with system-level integration complexity.
+Although the realization of large-scale ESPD architectures would require further advances in device integration and experimental implementation, the theoretical formulation and parameter studies presented here establish a concrete foundation for the framework and provide concrete guidance for future experimental efforts.
+In essence, this work establishes a conceptually complete and theoretically grounded framework for high-performance single-photon detection at room temperature.
+It demonstrates that intrinsic device-level limitations need not be overcome solely through component replacement, but can instead be effectively compensated by system-level architectural design.
+As such, the ESPD framework offers a viable pathway toward next-generation photon detection and provides a general methodology for rethinking performance limits in quantum technologies under practical constraints.
+Acknowledgements
+The author acknowledges the use of Gemini and ChatGPT (accessed December 2025) to, and only to, support language polishing, flow paragraphs, and correct grammar, to improve readability.
+All AI-generated suggestions were reviewed, revised, and approved by the authors, who take full responsibility for the accuracy and integrity of the work.
+All insight, results, methods, analysis, and experiments are ordinarily created and owned by the author.
+Code Availability
 All data/codes generated or used during this study are included in this published article or can be found at https://github.com/Hao-B-Shu/ESPD.
-Author Contributions
-Hao Shu (the unique author): study design, investigation, experiment, analysis, manuscript preparation.
-Declarations
-The author declares no competing interests.

### v5

@@
-Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature
+From Device to Dynamics: An Iterative Architectural Framework for High-Performance Single-Photon Detection at Room Temperature
 Abstract
-Photon detection, a fundamental quantum technology, is traditionally treated as a fixed device-level operation governed by intrinsic properties of single-photon detectors (SPDs).
-High-performance detection has therefore largely relied on superconducting technologies whose requirement for cryogenic operation imposes substantial infrastructure constraints, limiting scalable deployments.
-Here, the enhanced single-photon detection (ESPD) framework is presented as a system-level theoretical paradigm that shifts photon detection from device-centric optimization to an integrated quantum-information-processing task, by reformulating it as an iteratively enhanced process integrating state-preparation, controlled operations, projective measurements, and multi-copy analysis.
-ESPD enables systematic performance enhancement through architectural design rather than material modification, thereby circumventing superconducting components, allowing high-performance detection using exclusively room-temperature hardware.
-Numerical simulations grounded in physically motivated parameters indicate that the ESPD framework can upgrade a conventional room-temperature SPD to effective DE exceeding 93% and DCR below $10^{-9}$, which are comparable to state-of-the-art superconducting SPDs and can significantly relax the minimal tolerable channel transmission rate in quantum communications.
-While physical realization would require further component integration, this work establishes a rigorous theoretical framework for enhancing detection performance through architectural quantum-information principles, providing a general blueprint for transcending device-level constraints and guiding the development of next-generation room-temperature quantum technologies.
+Photon detection is a cornerstone of quantum technology, traditionally regarded as a static device-level operation constrained by the intrinsic physical properties of single-photon detectors (SPDs).
+Consequently, high-performance detection has been heavily reliant on superconducting technologies, whose requirement for cryogenic temperatures imposes significant infrastructure burdens and limits scalable deployment.
+To circumvent these constraints, we propose the Enhanced Single-Photon Detection (ESPD) framework, which shifts the photon-detection paradigm from device-centric optimization to an integrated quantum-information-processing (QIP) task.
+By incorporating state preparation, controlled operations, projective measurements, and multi-copy decision analysis, we establish a nonlinear dynamical model that reformulates detection as an iteratively enhanced process.
+This architecture enables systematic performance upgrades through structural design rather than material modification, allowing high-performance detection with exclusively room-temperature hardware.
+Through analytical approximations, Monte Carlo analysis, and numerical simulations, grounded in parameters derived from non-superconducting components, we show that the ESPD dynamics converge to a high-performance basin of attraction even when initialized by low-performance SPDs.
+Specifically, the framework can upgrade a conventional room-temperature SPD to achieve an effective detection efficiency (DE) exceeding 93% with a dark count rate (DCR) below $10^{-9}$, metrics comparable to state-of-the-art superconducting nanowire SPDs (SNSPDs).
+Such enhancements significantly lower the threshold for the required minimum channel transmission rate in quantum communication applications.
+While physical realization requires further component integration efforts, this work establishes a rigorous theoretical foundation for enhancing detection via architectural QIP principles.
+It provides not only a blueprint for next-generation room-temperature photon detection but also a general methodology for transcending device-level constraints in broader quantum technologies.
 keywords:
-Single-photon detection, System-level quantum architecture, Iterative quantum information processing, Room-temperature quantum technologies, Photonic integration
-Hao Shu*
+Single-photon detection, Room-temperature quantum technologies, Iterative enhanced quantum information processing, Nonlinear dynamics, Photonic integration
 […]
-{affiliations}
-Dr. Hao Shu
-Address: Sun Yat-Sen University, Guangzhou, China
-Email Address: Hao_B_Shu@163.com (Primary), shuh6@mail.sysu.edu.cn (Institutional)
+organization=Sun-Yat-Sen University,city=Shenzhen, country=China
 1 Introduction
-Single-photon detection plays a central role in quantum optics and serves as the bedrock for a wide range of quantum-information applications.
-Conventionally, the detection process is treated as a fixed device-level operation primarily determined by the material properties of single-photon detectors (SPDs).
-The primary performance metrics characterizing an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$.
-The former directly dictates the success probability of quantum operations conditioned on detection events, while the latter imposes a fundamental signal-to-noise constraint in high-attenuation scenarios.
-In terms of DE, high-performance detection is indispensable for retrieving meaningful outcomes in quantum experiments.
-In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [1, 2, 3].
+Photon detection is a foundational pillar of quantum optics, providing essential interfaces for nearly all quantum information applications.
+Traditionally, the detection process has been conceptualized as a static, device-level operation, whose performance is regarded as intrinsic to the single-photon detector (SPD) material properties.
+The utility of an SPD is primarily governed by two competing metrics: detection efficiency (DE, denoted by $\eta$), which represents the probability of detecting a non-vacuum incident photon, and the dark count rate (DCR, denoted by $d$), which represents the probability of returning an incorrect positive detection report in the absence of a photon.
+High DE is indispensable for obtaining meaningful outcomes in quantum experiments.
 In fundamental tests of quantum nonlocality, verifying violations of the Clauser-Horne-Shimony-Holt or Eberhard inequalities requires DE exceeding strict thresholds to close detection loopholes [1, 2, 3, 4, 5, 6, 7], such as approximately 67% [8] or 83% [9].
+In optical quantum computation, where successful events can be probabilistic and require the simultaneous detection of multiple photons, the overall success probability decreases exponentially with imperfect DE [10, 11, 12].
 In heralded single-photon sources, where one photon of an entangled pair signals the presence of the other, the achievable heralding rate is also strictly limited by DE [13, 14, 15, 16].
-With respect to DCR, a particularly stringent constraint arises in quantum communication, notably in quantum key distribution (QKD).
-The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are significantly lower than 50% [17, 18, 19, 20].
+Regarding DCR, a particularly stringent constraint arises in quantum communication, notably quantum key distribution (QKD).
+The security of practical QKD protocols relies on maintaining the quantum bit error rate (QBER) below protocol-dependent thresholds that are usually lower than 11% [17, 18, 19, 20, 21].
 However, optical communication channels inevitably induce exponential photon attenuation with distance [22, 23, 24, 25].
-When the effective signal rate, determined by channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts, which contribute a 50% error rate, causing the QBER to rise precipitously.
+When the effective signal rate, determined by exponentially decay channel transmission for a given DE, approaches the DCR floor of the SPDs, detection events become dominated by dark counts that contribute a 50% error rate, causing the QBER to rise precipitously.
 Consequently, the achievable secure distance of QKD systems is ultimately bounded by the DCR of the employed SPDs [18, 26, 27].
 Motivated by these critical demands, the development of high-performance SPDs has remained a primary research focus.
-While conventional semiconductor-based SPDs offer the advantage of non-cryonic operation, they typically suffer from an intrinsic trade-off between limited DE and relatively high DCR.
-For instance, a frequency upconversion SPD demonstrated in 2004 operated at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operated at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29].
-Such specifications often fall short of the requirements for advanced quantum protocols, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
-To meet these stringent requirements, substantial effort has shifted toward designing SPDs with superconducting technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33].
+While conventional semiconductor-based SPDs offer the advantage of non-cryogenic operation, they typically suffer from intrinsic limits on relatively low DE and high DCR.
+For instance, a frequency upconversion SPD reported in 2004 operates at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP SPD reported in 2017 operates at 223 K achieved a DE of about 27.5% with a DCR of $10^{-6}$ [29].
+Such specifications often fall short of advanced quantum task requirements, which frequently demand DE exceeding 90% and DCR below $10^{-7}$.
+To meet these stringent requirements, substantial efforts have gravitated toward superconducting SPD technologies, most notably superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33].
 These devices have successfully pushed DE beyond 90% while suppressing DCR by several orders of magnitude compared to semiconductor counterparts, thereby enabling demanding applications such as long-distance QKD [20, 34, 35, 36].
 Recent representative benchmarks of state-of-the-art superconducting SPDs are summarized in Table 1.
-Table 1: Summary of recent representative performance benchmarks for state-of-the-art superconducting SPDs.
-Year Researchers (et al.) Operation Temperature Detection Efficiency Dark Count Rate
+Table 1: Recent representative benchmarks of superconducting SPDs. OT: Operating Temperature; DE: Detection Efficiency; DCR: Dark Count Rate
+Year Researchers (et al.) OT DE DCR
 2020 P. Hu   [30] 2.10 K 95.0 % $0.5\times 10^{-5}$
@@
 These constraints present major barriers to scalable deployment, particularly in field applications, satellite payloads, and scenarios with strict size, weight, and power limitations where cryogenic infrastructure is impractical.
-To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, the enhanced single-photon detection (ESPD) framework is introduced, shifting the focus from device-centric optimization to a system-level architectural paradigm.
-Within this framework, single-photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on the material properties of a single SPD.
-Through structured system-level integration, the ESPD framework enables effective enhancement of detection performance using exclusively room-temperature components.
-Numerical simulations, grounded in physically motivated parameters drawn from commercially available devices and experimentally reported SPDs, indicate that, under an exclusively room-temperature component setting, the ESPD framework enables effective DE exceeding 93% together with effective DCR below $10^{-9}$, starting from a legacy SPD (for example, the one reported two decades ago with about 59% DE and $10^{-2}$ DCR [28]).
-These effective performance metrics are comparable to those of state-of-the-art superconducting SPDs and lead to a substantial relaxation of the minimal tolerable channel transmission rate in QKD.
-While physical implementation would require further system-level integration of quantum technologies, the ESPD framework provides a structured theoretical basis for achieving high-performance photon detection at room temperature.
-By explicitly analyzing technological feasibility and the trade-off between resource overhead and operating temperature, the ESPD framework offers a feasible technological roadmap and practical guidance for future experimental efforts.
-It serves as a theoretical blueprint for overcoming intrinsic device limitations, with the potential to unlock a wide range of quantum applications in environmentally constrained scenarios.
-2 Method and Result
-This section details the architectural structure of the ESPD framework and formulates the corresponding detection process.
-The ESPD paradigm is introduced in Section 2.1, with theoretical formulations established in Section 2.2.
-Analytical approximations are developed in Section 2.3, and numerical simulations are presented in Section 2.4.
-Finally, implications for QKD are discussed in Section 2.5.
+To address the limitations of room-temperature SPDs without resorting to cryogenic conditions, we introduce the enhanced single-photon detection (ESPD) framework, which shifts the focus from device-centric optimization to a system-level architectural paradigm with nonlinear dynamics.
+Within this framework, photon detection is reformulated as a composite quantum-information-processing task that integrates state preparation, controlled quantum operations, projective measurements, and multi-copy decision analysis, rather than relying solely on SPD material properties.
+By formulating this multi-stage enhancement approach as a non-linear dynamical system, we analyze the convergence range and steady-state performance of the ESPD structure, providing a novel perspective on hardware noise suppression through iterative convergence processing.
+Our analysis demonstrates that this framework can elevate the performance of conventional, room-temperature SPDs to levels comparable to, or even exceeding, those of cryogenic superconducting counterparts.
+Numerical simulations, grounded in physically motivated parameters from commercially available devices, indicate that the iterative mapping within the ESPD framework exhibits strong convergence toward high-performance regimes.
+Monte Carlo analysis reveals that, provided the initial SPD parameters fall within a broad basin of attraction (i.e., not starting with an extremely poor SPD), the evolved DE can surpass 90% alongside ultra-low DCR.
+For instance, starting from a legacy SPD (e.g., DE $\approx 59\%$ and DCR $\approx 10^{-2}$ [28]), the system converges to an effective DE exceeding 93% and an effective DCR below $10^{-9}$.
+Such performance rivals or even outperforms state-of-the-art superconducting SPDs, substantially relaxing the minimal tolerable channel transmission rate in QKD, and is achieved under room-temperature settings, circumventing the need for cryogenic infrastructure.
+While physical implementation will require further system-level integration of quantum technologies, the ESPD framework provides a structured nonlinear dynamical formulation for room-temperature high-performance photon detection.
+Furthermore, we offer a practical roadmap to guide future experimental efforts by explicitly discussing technological feasibility and analyzing the trade-offs between resource overhead and operating temperature.
+Ultimately, the design principles and perspective shift offered by the ESPD framework serve as a theoretical blueprint for overcoming intrinsic device limitations through dynamical process design, which can boost the development of a broader range of quantum devices beyond photon detection and unlock diverse high-performance quantum applications at room temperature.
+2 Methods and Results
+This section details the ESPD architecture and formulates the corresponding detection dynamics.
+The ESPD paradigm is introduced in Section 2.1, with rigorous theoretical formulations established in Section 2.2.
+Analytical approximations are developed in Section 2.3, followed by (Monte Carlo) stability analysis and numerical simulations in Section 2.4.
+Finally, the implications for QKD applications are discussed in Section 2.5.
 2.1 Paradigm of the Enhanced Single-Photon Detection
-The ESPD architecture is formulated as a level-by-level framework in which the effective detection performance of a given SPD is enhanced through successive processing stages.
-Here, the enhancement is quantified in terms of DE and DCR, defined respectively as the probability of generating a positive detection event given a non-vacuum (single-photon) input and the probability of a positive report given a vacuum input.
-Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with effective DE $\eta_{s}$ and effective DCR $d_{s}$.
-The iteration begins with an existing SPD, denoted as $\text{ESPD}_{0}$, characterized by the initial values $\eta_{0}=\eta$ and $d_{0}=d$.
-With qubit logic, the schematic of the ESPD paradigm is illustrated in Fig. 1.
-Figure 1: Schematic representation of the recursive ESPD architecture.
-The $(s+1)$-th level detector ($\text{ESPD}_{s+1}$) is constructed by integrating multiple copies of the $s$-th level detector ($\text{ESPD}_{s}$) with controlled gates.
-Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective measurements via basis $\{|0\rangle,|1\rangle\}$, and filters represent that the $|0\rangle$ path is aborted.
+The ESPD architecture is formulated as a recursive, level-by-level framework in which baseline SPD performance is systematically enhanced through successive processing stages.
+Within this paradigm, the performance is characterized by the effective DE ($\eta_{s}$) and DCR ($d_{s}$) at level $s$.
+Recall that DE is the probability of generating a positive detection event given a non-vacuum (single-photon) input, and DCR is the probability of generating an (incorrect) positive report given a vacuum input.
+Let $\text{ESPD}_{s}$ denote the established detector at level $s$, characterized by $(\eta_{s},d_{s})$.
+The evolution begins with an existing SPD, denoted as $\text{ESPD}_{0}$, with initial values $\eta_{0}=\eta$ and $d_{0}=d$, the performance of the physical baseline SPD.
+The schematic of this recursive logic is illustrated in Fig.
+1.
+Figure 1: Schematic representation of recursive ESPD architecture.
+The $(s+1)$-th level detector ($\text{ESPD}_{s+1}$) is constructed by integrating multiple copies of the $s$-th level detector ($\text{ESPD}_{s}$) with state preparation, controlled gates, and projective measurements.
+Here, $C(F_{i})$ are controlled gates, SP are state-preparation operations, measurements are projective in the basis $\{|0\rangle,|1\rangle\}$, and filters indicate the abortion of the $|0\rangle$ path.
 The construction of the $(s+1)$-th level detector, $\text{ESPD}_{s+1}$, is defined as follows.
-As a representative example, an implementation may be considered in which the degree of freedom (DOF) is photon polarization and the controlled operations correspond to controlled-NOT (C-NOT) gates.
+As a representative example, an implementation can be considered with the degree of freedom (DOF) using photon polarization and the controlled operations using controlled-NOT (C-NOT) gates.
   1. 1.
-    DOF Selection: A manipulable DOF of the incoming photon is chosen (for example, polarization).
+    DOF Selection: Select a manipulable DOF of the incoming photon (for example, polarization).
 The computational basis for this DOF is denoted by $\{|0\rangle,|1\rangle\}$.
-The signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit.
+Hence, the signal is modeled as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle$ and $|1\rangle$ denote orthonormal states of a non-vacuum photon [^1] If the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents a non-vacuum state, reducing the description to qubit.
 In practical architectures, an internal DOF such as polarization, distinct from photon existence, is often more convenient..
   2. 2.
-    State Preparation: A state-preparation operation is applied to initialize the selected DOF to $|1\rangle$.
+    State Preparation: Apply the state-preparation operation, initializing the selected DOF to $|1\rangle$.
 As a result, a non-vacuum signal is mapped to $|1\rangle$, while a vacuum input remains in state $|\emptyset\rangle$.
   3. 3.
-    Controlled Operations: A set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), are applied to couple the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
+    Controlled Operations: Operate a set of $n_{s+1}$ controlled operations, denoted as $C(F_{i})$ for $i=1,\dots,n_{s+1}$ (typically C-NOT gates), coupling the signal path with $n_{s+1}$ auxiliary photons, each initialized in state $|0\rangle$.
   4. 4.
-    Auxiliary Measurement: Projective measurements are performed on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$.
-Outcome paths corresponding to $|0\rangle$ are discarded, and only the $|1\rangle$ outcomes are retained for subsequent detection.
+    Auxiliary Measurement: Perform projective measurements on the auxiliary outputs in basis $\{|0\rangle,|1\rangle\}$.
+Outcome paths corresponding to $|0\rangle$ are discarded; only the $|1\rangle$ outcomes are retained for subsequent detections.
   5. 5.
-    Detection: Detection is carried out on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
+    Detection: Carried out detections on all remaining paths, including both the signal path and the auxiliary paths, using $s$-th level detectors ($\text{ESPD}_{s}$).
   6. 6.
-    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path.
+    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detection events occur among the $n_{s+1}+1$ detectors, consisting of $n_{s+1}$ auxiliary detectors and one detector on the signal path.
 Otherwise, a negative report is returned.
@@
 In this case, the decision logic is naturally reformulated in terms of the fraction of positive detection outcomes among the effective detections exceeding a predefined threshold..
-Physical intuition: From a detection perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal.
-In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ effective signal instances through the controlled modules, which are independently detected by $n_{s+1}$ SPDs.
-A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered.
-Under this idealized assumption, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$.
-Furthermore, the effective DCR is suppressed, since a false positive outcome requires dark-count events to occur in at least $k_{s+1}$ detectors, leading to an approximate scaling $d_{s}^{k_{s+1}}$.
-This physical intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal.
-A rigorous formulation of this behavior, accounting for non-ideal devices and error propagation, is developed in Section 2.2.
+Physical intuition: From a functional perspective, each level of the ESPD framework increases the number of effective detection opportunities available to an input signal.
+In an idealized limit where all components except the SPDs are assumed to be perfect, the $(s+1)$-th level produces $n_{s+1}$ auxiliary signal instances identical to the input one through the controlled modules, which are independently detected by the $n_{s+1}$ ESPDs.
+A non-vacuum outcome is reported if at least $k_{s+1}$ positive detection events are registered among the $n_{s+1}+1$ detections.
+Therefore, the effective DE increases with decreasing threshold $k_{s+1}$; for instance, when $k_{s+1}=1$, the effective DE is given by $1-(1-\eta_{s})^{n_{s+1}}$.
+Furthermore, the effective DCR is suppressed because a false positive outcome requires at least $k_{s+1}$ independent dark count events, leading to an approximate probability of $d_{s}^{k_{s+1}}$.
+This architectural logic transforms detection from a single-shot device event into a more robust collective process, where multi-copy analysis can suppress the errors introduced by SPD imperfections.
+This intuition extends beyond the idealized limit when imperfections in the auxiliary components introduce errors that remain small compared to the activation probability of a genuine input signal.
+A rigorous treatment of these dynamics, accounting for non-ideal devices and error propagation, follows in Section 2.2.
 2.2 Theoretical Formulations
-The rigorous theoretical formulation of the ESPD framework is developed in this subsection.
-The relevant notations and assumptions are introduced first, followed by the exact analytical expressions and the approximation analysis.
+This subsection develops the rigorous theoretical foundation of the ESPD framework.
+We first define the governing parameters and underlying assumptions, followed by deriving the analytical expressions of the ESPD dynamics.
 2.2.1 Notations and Basic Assumptions
-The parameters governing the system behavior are defined as follows.
-As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively.
-Although $\eta_{s}$ may depend on the photon number distribution, it is approximated as a unified value for the given system for simplicity.
-The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
-For convenience, all controlled operations $C(F_{i})$ are assumed to be identical and are denoted simply as $C(F)$.
-They are implemented in a feedforward controlled configuration, such that $n_{s+1}$ represents both the number of effective modules and the total number of physical modules required at level $s+1$.
+The evolution of the system is governed by a set of parameters characterizing both detection components and coupling modules.
+As defined previously, $\eta_{s}$ and $d_{s}$ represent the effective DE and DCR of the $\text{ESPD}_{s}$ unit, respectively.
+For analytical tractability, $\eta_{s}$ is treated as a unified value under a given photon-number distribution in the system.
+The architectural configuration at each level is determined by the number of effective controlled operations $n_{s}$ and the decision threshold $k_{s}$.
+We assume that all controlled operations $C(F_{i})$ are identical (denoted as $C(F)$) and implemented in a feed-forward configuration, such that $n_{s}$ represents both the number of effective modules and the total number of physical modules required at level $s$.
 However, the generalized ESPD paradigm remains valid for post-selection controlled gates (see Section 3.2.2 for a brief discussion).
-Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$).
+Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$), accounting for incidental losses.
 Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared).
-Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively.
-Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection).
-The parameters $p$, $P$, and $Q$ are approximately assumed to be invariant across all $C(F)$ module operations.
+We further define $P$ and $Q$ as the conditional probabilities of obtaining a non-vacuum output on the auxiliary (controlled) path before detection, given a non-vacuum or vacuum input to the module, respectively.
+Thus, $Q$ encapsulates the intrinsic error rate of the auxiliary path attributable to state preparation and measurement (SPAM), not including the final detection, in the absence of $C(F)$.
 It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module.
 These parameters are defined at the component level and remain invariant with respect to the number of cascaded modules.
-The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
+Their cumulative impact on the system’s evolutionary trajectory is fully captured by the recursive relations derived in Section 2.2.3.
 2.2.2 Parameter Approximation Assumptions
-The assumptions below are employed solely to derive analytical approximations in Section 2.3, as well as the approximation part in Eq.
-(1) to (3).
-They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work.
-Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section 3.1.
+While the general formulation remains exact, the following assumptions, which are physically and practically motivated, are introduced to facilitate the analytical approximations in Section 2.3.
+Such assumptions are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations and will be discussed in detail in Section 3.2.
 In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$.
@@
 Finally, considering practical implementation costs, $n_{s}$ and $k_{s}$ are taken to be moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
-2.2.3 Formulations
-Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum and vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector.
-They can be expressed as:
+2.2.3 Evolutionary Equation Formulations
+To describe the state transition from level $s$ to $s+1$, we first evaluate the reporting probabilities on the auxiliary and signal paths.
+Let $P_{s+1}$ and $Q_{s+1}$ be the probabilities of obtaining a positive report on an auxiliary path using $\text{ESPD}_{s}$ as the detector, given a non-vacuum and vacuum input to the $C(F)$ module, respectively:
 $P_{s+1}=P[\eta_{s}+(1-\eta_{s})d_{s}]+(1-P)d_{s}=P\eta_{s}(1-d_{s})+d_{s}\approx P\eta_{s}$ (1)
@@
 Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum and vacuum output from the final controlled gate, respectively, then:
-$P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\ Q^{\prime}_{s+1}=d_{s}$ (3)
-The DE and DCR for the $(s+1)$-th level are derived as follows.
+$P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\quad Q^{\prime}_{s+1}=d_{s}$ (3)
+The performance of $\text{ESPD}_{s+1}$ is determined by the collective response of $n_{s+1}+1$ detectors.
 For the DE calculation, a non-vacuum input is assumed, and the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections is evaluated.
 If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
-$P_{s+1,k_{s+1},i}=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$ (4)
+$P_{s+1,k_{s+1},i}$ (4)
+$=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
 $+Q^{\prime}_{s+1}\sum_{j_{1}+j_{2}=k_{s+1}-1}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
@@
 If the photon survives all $C(F)$ operations, the probability of a final positive report is:
-$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
-$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
-By combining the above contributions, the DE of $\text{ESPD}_{s+1}$ is given by:
+$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
+$\qquad\qquad\qquad\ +(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
+$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
+$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
+The total effective DE at level $s+1$ is then the weighted sum:
 $\eta_{s+1}=p^{n_{s+1}}P_{s+1,k_{s+1}}+\sum_{i=1}^{n_{s+1}}p^{i-1}(1-p)P_{s+1,k_{s+1},i}$ (6)
 For the DCR calculation, a vacuum input is assumed. The probability of a false positive is given by:
-$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
-Collectively, the DE and DCR evolution is governed by the non-linear time-variant dynamic system:
-$(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\ (\eta_{0},d_{0})=(\eta,d)$ (8)
-where the map $G_{s+1}$ is defined by Eq. (1) to (7).
-2.3 Analytical Approximations and Stability Analysis
+$d_{s+1}=(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}$ $\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
+$+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
+In summary, the transition of performance metrics follows a non-linear dynamical system:
+$(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\quad(\eta_{0},d_{0})=(\eta,d)$ (8)
+where the map $G_{s+1}$ is defined by the recursive operators in Eqs. (1) to (7).
+The ESPD architecture thus induces a nonlinear dynamics whose fixed points, stability, and basins of attraction determine the asymptotic detection performance.
+Particularly, the fixed point $(\eta_{\infty},d_{\infty})$ represents the fundamental physical limit of the ESPD framework under given architectural parameters, while the basins of attraction define the range of hardware-level imperfections that can be effectively mitigated to achieve such enhancement.
+2.3 Analytical Approximations Analysis
 While Eq.
-(8) describes the exact dynamics, an analytical solution is generally intractable.
-However, by employing the approximations in Eq.
-(1) to (3) together with the assumptions in Section 2.2.2, simplified bounds can be derived to elucidate the system behavior.
-The evolution of the DCR is considered first. Eq. (7) can be reformulated as:
-$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (9)
+(8) describes the exact dynamics of the ESPD framework, an explicit analytical solution for the iterated map is generally intractable.
+However, by employing the approximations in Eqs.
+(1) to (3) alongside the physically-motivated assumptions in Section 2.2.2, we can derive simplified bounds to elucidate the system’s evolutionary behavior.
+Consider the evolution of DCR first. Eq. (7) can be reformulated as:
+$d_{s+1}=Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1$- $Q_{s+1})^{n_{s+1}-k_{s+1}+1}$ (9)
+$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
+To analyze the monotonicity of this transition, we introduce the following proposition.
 Proposition 1.
@@
 Proof: For $x\in[0,\frac{k-1}{n}]$:
-$\frac{d}{dx}f(x)=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$ (11)
-$+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
+$\frac{d}{dx}f(x)$ (11)
+$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$
+$\qquad\qquad\qquad\quad+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
 $=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}(k-1-nx)+\sum_{j\geq k}\binom{n}{j}x^{j+1-k}(1-x)^{n-j-1}(j-nx)]$
@@
 $\blacksquare$
-By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, the following bound is obtained:
-$d_{s+1}\leq$ $d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$ (12)
-which is notably independent of $\eta_{s}$.
-Given $Q_{s},d_{s}\ll 1$ and moderate values of $n_{s+1}$ and $k_{s+1}$, it is assumed that $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$.
+By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, we obtain the following bound for the DCR evolution:
+$d_{s+1}\leq d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})$ ${}^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}$ (12)
+$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$
+Notably, this bound is independent of the DE $\eta_{s}$.
+Given $Q,d_{s}\ll 1$ and moderate values for the architectural parameters $(n,k)$, we can assume $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$.
 This yields the approximation:
-$d_{s+1}\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$ (13)
-This provides an estimation for the DCR at level $s+1$.
+$d_{s+1}$ $\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}$ (13)
+$=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$
+This expression provides a practical estimation for the DCR suppression at the iterative level $s+1$.
 Similarly, the DE can be lower-bounded by:
-$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$ (14)
-$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
-$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
-which is independent of $d_{s}$. Therefore, the condition for DE enhancement at the subsequent level is determined by solving the inequality
+$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (14)
+$\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
+$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
+$\qquad\qquad\qquad\qquad\qquad\qquad\ +\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
+$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}$
+$\qquad\qquad\qquad\qquad\qquad\qquad\ +\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
+This bound is independent of $d_{s}$. Hence, the condition for DE enhancement at the subsequent level (namely $\eta_{s+1}>\eta_{s}$) is determined by the inequality:
 $p^{n}(Px)^{k-1}[x\binom{n}{k-1}(1-Px)^{n-k+1}+\sum_{j\geq k}\binom{n}{j}(Px)^{j-k+1}(1-Px)^{n-j}]-x>0$ (15)
-If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
-2.4 Numerical Simulations
-While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design.
-In this section, numerical simulations of the ESPD paradigm are presented, indicating that substantial performance gains can be achieved even with moderate initial hardware specifications.
-For example, a three-level ESPD can upgrade a semiconductor SPD to effective performance metrics with DE $93\%$ and DCR $10^{-9}$, comparable to those of state-of-the-art SNSPDs without reliance on superconducting materials.
-Simulation Parameters: The performance of the ESPD scheme is evaluated using two distinct baseline SPDs.
-The first represents an older one, characterized by parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29].
-Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section 3.1 for parameter discussions).
-Supplementary simulations with degraded parameter settings are also performed to evaluate robustness in Section 3.1.
-Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig.
-2.
-The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs.
-Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
-Table 2: Performance of an ESPD initialized with a SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
-Para 1 Para 2 Para 3
+If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of the $\text{ESPD}_{s}$ will (approximately) converge to a stable fixed point corresponding to a root of the left-hand side of Eq.
+(15).
+2.4 Monte Carlo Steady-State Analysis and Numerical Simulations
+While obtaining an exact analytical solution for the non-linear dynamic system governing the ESPD in Eq.
+(8) is mathematically complex, numerical methods provide a robust approach for stability analysis and practical design.
+In this section, we present the Monte Carlo analysis to characterize the system’s evolutionary behavior, including the location of stable fixed points and their basins of attraction, and provide numerical simulations under practically motivated device parameters.
+The results indicate that substantial performance gains can be achieved even with moderate initial hardware specifications.
+For instance, a three-level ESPD can elevate a conventional semiconductor SPD to effective performance metrics (DE $93\%$ and DCR $10^{-9}$), matching state-of-the-art SNSPDs, but without requiring cryogenic infrastructure.
+Simulation Parameters: The parameter settings are physically motivated by commercial specifications or experimental reports.
+The controlled-gate parameters are set to $p=0.98$ and $P=0.97$, values recently demonstrated in optical platforms and feedforward C-NOT gate implementations on photonic platforms [38, 39, 40, 41, 42].
+For the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation [43, 42].
+Detailed discussions on parameters and robustness under degraded settings are provided in Section 3.2.
+[…]
+Performance Evolution on Legacy SPD: We specifically examine the evolutionary trajectory of the ESPD initialized with a legacy SPD from [28] (DE $\approx 59\%$, DCR $\approx 10^{-2}$).
+Table 2 details the step-by-step evolutions, with visual representations in Fig.
+5.
+These results illustrate that the ESPD scheme significantly transcends the intrinsic limitations of conventional SPDs.
+Notably, within only three to four levels of iterative enhancement, the effective metrics (DE $93\%$, DCR $10^{-9}$) become comparable to advanced superconducting SPDs, consistent with the stability predicted by our steady-state analysis.
+(a) Steady-state DE $\eta_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
+(b) Steady-state DCR $d_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
+Figure 2: Monte Carlo steady-state analysis of DE and DCR of the ESPD framework for $(n_{s},k_{s})=(8,4)$ at each level $s$.
+(a) Steady-state DE $\eta_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
+(b) Steady-state DCR $d_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
+Figure 3: Monte Carlo steady-state analysis of DE and DCR of the ESPD framework for $(n_{s},k_{s})=(6,3)$ at each level $s$.
+(a) Steady-state DE $\eta_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
+(b) Steady-state DCR $d_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
+Figure 4: Monte Carlo steady-state analysis of DE and DCR of the ESPD framework for $(n_{s},k_{s})=(4,2)$ at each level $s$.
+Table 2: Performance of an ESPD initialized with an SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
+Para 1: $(n,k)=(4,2)$ Para 2: $(n,k)=(6,3)$ Para 3: $(n,k)=(8,4)$
 Level $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$)) $(n,k)$ (DE ($\eta$), DCR ($d$))
 […]
-1 (4,1) $(97.4\%,\ 5.3\times 10^{-2})$ (6,1) $(98.4\%,\ 7.5\times 10^{-2})$ (8,1) $(98.6\%,\ 9.5\times 10^{-2})$
+1 (4,2) $(97.4\%,\ 5.3\times 10^{-2})$ (6,3) $(98.4\%,\ 7.5\times 10^{-2})$ (8,4) $(98.6\%,\ 9.5\times 10^{-2})$
 2 (4,2) $(98.2\%,\ 2.7\times 10^{-2})$ (6,3) $(96.6\%,\ 1.2\times 10^{-2})$ (8,4) $(95.1\%,\ 7.4\times 10^{-3})$
@@
 8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-Table 3: Performance of an ESPD initialized with a SPD from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 5: Performance of the ESPD initialized with SPDs from [28] with $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
+2.5 Implications for QKD
+As a representative application, we analyze the implications of the ESPD framework applied to QKD.
+In a QKD protocol characterized by a secure error threshold $e_{th}$ and a baseline error rate $e_{C}$ (excluding detector-induced errors), the minimal tolerable channel transmission rate $\gamma$, which defines the operational limit of a link, is conventionally approximated by[27]:
+$\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
+Since the ESPD paradigm can substantially suppress effective DCR $d$ while simultaneously enhancing the DE $\eta$, it enables a dramatic reduction in $\gamma$.
+This suppression of the minimal transmission threshold directly translates into an extended communication distance and increased robustness against channel loss [^3] The channel transmission rate $t$ is conventionally calculated by $t=10^{-\frac{\alpha l}{10}}$, where $\alpha$ is a fixed parameter mainly determined by channel characters and $l$ is the communication distance..
+Furthermore, the enhancement of $\eta$ not only lowers the threshold $\gamma$ but also scales the achievable secret key rate at any fixed channel transmission, significantly improving the system throughput.
+Compared to existing alternative strategies, such as the empty-signal-detection paradigm [27], the ESPD framework offers a distinct functional advantage.
+While the previous method also reduces the channel transmission threshold, it introduces additional sifting consumption that inherently degrades the efficiency.
+In contrast, the ESPD framework achieves performance gains by directly improving DE and reducing DCR, without the need for auxiliary sifting.
+Consequently, the ESPD paradigm provides a more efficient approach for improving QKD performance under practical constraints.
+3 Discussion
+In this section, we discuss the technological feasibility and practical robustness of the ESPD framework, as well as its limitations and directions for future work.
+3.1 Effectiveness when the Initial SPD is Outside the Convergent Range
+As evidenced by the Monte Carlo analysis, the steady-state performance of the ESPD is largely independent of the specific characteristics of the initial SPD, provided that it lies within a convergence basin.
+However, although the convergence interval is relaxed, some traditional SPDs might still fall outside this range.
+For example, the InGaAs/InP SPD designed in 2017 with a relatively lower DE ($\approx 27.5\%$) and lower DCR ($\approx 10^{-6}$)[29] is outside the range in which the ESPD dynamics converge to a high-performance fixed point when a uniform configuration of $(n_{s},k_{s})=(6,3)$ or $(n_{s},k_{s})=(8,4)$ is shared by each level.
+Nevertheless, such SPDs can still be significantly enhanced by the ESPD framework through an appropriate selection of the first-level configuration parameters $(n_{1},k_{1})$.
+This strategic parameter choice allows states of the dynamical system described by Eq.
+(8) to be driven into the basin of attraction at the first step.
+For instance, selecting $(n_{1},k_{1})=(6,1)$ and $(n_{1},k_{1})=(8,1)$ results in $(\eta_{1},d_{1})=(85.5\%,3.4\times 10^{-3})$ and $(\eta_{1},d_{1})=(90.0\%,4.5\times 10^{-3})$, respectively, starting from the aforementioned SPD from [29], which successfully falls within the convergence basin for subsequent levels.
+Indeed, while a trade-off exists between DE and DCR, DCR of the ESPD can be drastically reduced by increasing $k$, and DE can be significantly enhanced by selecting a small $k$.
+Therefore, by employing one or two levels of the ESPD paradigm with suitable configuration parameters as a buffer, the detector’s performance can be adjusted into a desirable operational window, as long as the initial SPD is not prohibitively low.
+Numerical simulations for an initial SPD with $(\eta_{0},d_{0})=(27.5\%,1.0\times 10^{-6})$ are provided in Table 3 and visualized in Fig.6, suggesting that SPDs with either low DE and low DCR [29], or moderate DE and high DCR [28], can serve as viable initializations.
+Table 3: Performance of ESPD initialized with an SPD from [29], with initial parameters $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.
 Para 1 Para 2 Para 3
@@
 8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
-(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
-(d) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$. Initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
-Figure 2: Performance of ESPD initialized with SPDs from [28] ((a), (b)) and [29] ((c), (d)), respectively.
-2.5 Implications for QKD
-As a representative application scenario, the implications of the ESPD framework for QKD are analyzed.
-For a QKD protocol characterized by a secure error threshold $e_{th}$ and a total error rate $e_{C}$ excluding detector contributions, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by[27]:
-$\gamma=\frac{(1-2e_{th})d}{\eta[e_{th}-e_{C}+d(1-2e)]}\approx\frac{1-2e_{th}}{e_{th}-e_{C}}\frac{d}{\eta}$ (16)
-Since the ESPD paradigm can substantially suppress the effective DCR $d$, the minimal tolerable channel transmission rate $\gamma$ is proportionately reduced.
-Furthermore, the simultaneous enhancement of the DE $\eta$ not only reduces $\gamma$ but also directly improves the achievable key rate at a fixed channel transmission rate.
-Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD framework achieves this improvement without introducing additional sifting procedures that reduce the effective sifting efficiency.
-Here, QKD performance can be enhanced through the simultaneous suppression of the effective DCR and enhancement of the effective DE, without the need for auxiliary sifting operations.
-In this sense, the ESPD paradigm provides a more efficient system-level approach for improving QKD performance under practical constraints.
-3 Discussion
-In the absence of an experimental demonstration, we focus on the technological feasibility and practical robustness of the ESPD framework, and discuss its limitations as well as directions for future work.
-3.1 Technological Feasibility and Performance under Degraded Parameters
-The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs.
-The device requirements are discussed as follows.
-3.1.1 State Preparations and Measurements
-The flexibility of the ESPD architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one.
-In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be achieved at extremely low levels.
-Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
+(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
+Figure 6: Performance of ESPD initialized with an SPD from [29] with initial parameters: $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$, respectively.
+3.2 Technological Feasibility and Robustness under Degraded Parameters
+The primary technological advantage of the ESPD paradigm lies in its exclusive reliance on accessible, room-temperature components, effectively circumventing the stringent cryogenic requirements inherent to superconducting SPDs.
+The technological feasibility of components is discussed as follows.
+3.2.1 State Preparation and Measurement
+The architectural flexibility of the ESPD framework allows for various DOFs to be employed, enabling the selection of an easily manipulated DOF based on specific platforms.
+In practical implementations, this facilitates high-fidelity SPAM, ensuring that the error parameter $Q$ can be maintained at a low level.
+Specifically, $Q$ can be suppressed to the $10^{-4}$ level using standard optical setups and commercial platforms [43, 42], guaranteeing $Q\ll 1$.
 3.2.2 Controlled Gates
 The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
-Table 4: Performance of an ESPD initialized with a SPD from [28], with $p$ varied from 0.80 to 0.96.
-Simulations are performed for $(n,k)=(5,2)$ at each level, and all other settings remain consistent with those described in Section 2.4.
+High transmission efficiencies, for example, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [42], are achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
+It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, remains invariant under an increasing number of cascaded modules or ESPD levels.
+The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
+Table 4: Performance of an ESPD initialized with an SPD from [28], with $p$ varied from 0.80 to 0.96.
+Simulations are performed for $(n,k)=(5,2)$ at each level $L$, and all other settings remain consistent with those described in Section 2.4.
 pp=0.80 pp=0.84 pp=0.88 pp=0.92 pp=0.96
-Level (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$))
+$L$ (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$)) (DE ($\eta$), DCR ($d$))
 0 $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$ $(59.0\%,1.0\times 10^{-2})$
@@
 (b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
-Figure 3: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
-A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity.
-It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module and, by definition, does not decrease with an increasing number of cascaded modules or ESPD levels.
-The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
+Figure 7: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$ with degraded pp. All other settings remain consistent with those described in Section 2.4.
 However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations.
-While the ESPD paradigm remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded.
+While the ESPD paradigm remains valid for such probabilistic operations (with slight modifications in analysis), the overall system efficiency is necessarily degraded.
 Specifically, if the post-selection efficiency of the controlled gate is $\frac{1}{N}$, the effective average transmission rate from one module to the next is reduced to $p^{N}$.
 Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate.
-Table 4 and Fig.3 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.).
+Table 4 and Fig.7 indicate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss accumulated across approximately 10 controlled modules with 0.1 dB attenuation on each.).
 This result shows robustness, indicating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 42].
@@
 Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity should account for the correct operation rate of both the control and controlled paths.
-Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38].
-Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryonic components[39, 40, 41, 42].
-However, it is also worth testing the effectiveness of the paradigm under degraded parameters to indicate its suitability for standard fabrication processes with lower technology requirements.
-Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), with visual representation in Fig.4.
+Contemporary room-temperature implementations for C-NOT demonstrate high performance, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 42].
+Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% at non-cryogenic components[38, 39, 40, 41].
+However, it is also worth further testing the paradigm’s resilience under degraded parameters.
+Here, we also provide simulation results for two degraded scenarios: $P=0.80$ (see Table 5), and $P=0.40$ (see Table 6), whose visual representation is shown in Fig.8 and 9.
 All other settings remain consistent with those described in Section 2.4.
-The numerical results indicate the robustness of the scheme: Comparison of Table 5 to Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$.
-This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with existing room-temperature quantum technologies.
-Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
+The numerical results indicate the robustness of the scheme: Comparison of Table 5 with Table 2 supports that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded, such that $P=0.40$.
+This robustness demonstrates that the technological requirements of the ESPD method are fully compatible with and near-term room-temperature quantum fabrication processes.
+Table 5: Performance evolution of the ESPD initialized with an SPD from [28], and moderate controlled gate fidelity, $P=0.80$.
 Para 1 Para 2 Para 3
@@
 8 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
-Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under a low controlled gate fidelity, $P=0.40$.
+(a) DE $\eta_{s}$ vs. level $s$.
+(b) DCR $d_{s}$ vs. level $s$.
+Figure 8: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
+Table 6: Performance evolution of the ESPD initialized with an SPD from [28], under a low controlled gate fidelity, $P=0.40$.
 Para 1 Para 2 Para 3
@@
 8 (6,2) $(75.4\%,\ 3.7\times 10^{-5})$ (7,2) $(83.9\%,\ 6.4\times 10^{-5})$ (8,2) $(87.7\%,\ 9.6\times 10^{-5})$
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
-(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.80$.
-(c) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$, with $P=0.40$.
-(d) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$, with $P=0.40$.
-Figure 4: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$ ((a), (b)) and $P=0.40$ ((c), (d)), respectively.
-3.1.3 Choice of initialized SPDs
-Notably, the requirements imposed on the SPD used for initialization are minimal, which further lowers the practical barrier to implementation.
-As evidenced by the comparison between Tables 2 and 3, the stable performance of the ESPD is largely independent of the specific characteristics of the initialized SPD, provided that it lies within a certain convergence interval.
-The dynamic system described by Eq.(8) exhibits convergence to a stable operational point within this interval, irrespective of the initial point.
-Moreover, while a trade-off exists between DE and DCR, the DCR of the ESPD can be drastically reduced by increasing $k$, and the DE can be significantly enhanced by selecting a small $k$.
-Therefore, one or two levels of the ESPD paradigm, using suitable $n$ and $k$, can readily adjust the detector’s performance to a desirable operational window, provided the initial SPD is not severely deficient.
-Indeed, numerical simulations suggest that both SPDs with low DE and low DCR[29], as well as those with moderate DE and relatively high DCR[28], can serve as viable initializations.
-In summary, the ESPD scheme requires only a conventional room-temperature SPD for initialization, which is routinely achievable with existing technology.
-For instance, the SPD developed two decades ago [28] is sufficient.
+(a) DE $\eta_{s}$ vs. level $s$.
+(b) DCR $d_{s}$ vs. level $s$.
+Figure 9: Performance for ESPD with initialized parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
 3.2.3 Auxiliary Sources
-Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single photon sources.
-The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, regardless of the exact number of photons or the presence of unwanted $|0\rangle$.
-Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required $|1\rangle$ [^3] Since the employment of the filter block, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects.
+Although nearly perfect sources can be obtained at room temperature [53], indicating that auxiliary sources will not be a barrier, implementing the ESPD scheme may not require single-photon sources.
+The key parameter $P$ represents the detection probability of the state $|1\rangle$ from the auxiliary path with an ideal SPD when the control signal input is non-vacuum, which is a global parameter regardless of the exact number of photons or the presence of unwanted $|0\rangle$.
+Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more detection opportunities for the required state $|1\rangle$ [^4] Since the architecture incorporates a filter block, any additional $|0\rangle$ components will be filtered out.
+Therefore, for a non-vacuum input of a controlled module, if the auxiliary multi-photon signal results in more state $|1\rangle$, then a positive report has a higher chance to be produced, which provides benefits since it enlarges $P$, while, if the auxiliary multi-photon signal adds state $|0\rangle$, then they will be filtered out, thus providing no effects.
 Similarly, for a vacuum input of a controlled module, only $|0\rangle$ is added in auxiliary paths, which will be filtered out and thereby have no effect even if the source is multiple-photon..
 However, a rigorous formulation of this effect, including its impact on the controlled operations, remains an open problem and is left for future investigation.
-3.2 Limitations and Future Works
+3.3 Limitations and Future Work
 Despite the significance of the ESPD framework, some limitations remain and warrant future investigation.
@@
 The theoretical framework established in Section 2.2 remains an area for deeper exploration.
-While the numerical designs presented in Section 2.4 are sufficient for guiding physical implementation, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest.
-Future work should focus on rigorously deducing the exact convergence conditions, calculating the precise steady-state point, and identifying the minimum initial performance criteria required.
-Given the inherent mathematical difficulty of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
-3.2.2 Resource Overhead vs. Environmental Simplicity
-While the parallelized nature of the ESPD’s detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that scales with the number of levels.
+While the Monte Carlo analysis and numerical simulations presented in Section 2.4 are sufficient for convergence demonstration and physical implementation guidance, a complete theoretical study of the dynamic system in Eq.(8) remains of fundamental academic interest.
+Future work should focus on rigorously deducing the exact convergence conditions and calculating the precise steady-state point from a more theoretical perspective.
+Given the inherent complexity of this non-linear and time-variant system, and the fact that numerical designs already suffice for practical implementations, such specialized theoretical investigations are deferred to future work.
+3.3.2 Resource Overhead vs. Optional Simplicity
+While the parallelized nature of the ESPD detection stages theoretically preserves efficiency, the recursive architecture introduces a resource overhead that increases exponentially with the number of levels.
 Specifically, an $L$-level ESPD utilizing $n$ auxiliary signals per level requires a total of $(n+1)^{L}$ base detections.
-Although effective performance is typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling significantly expands the physical footprint and component count compared to a single SNSPD.
-For instance, a three-level ESPD scheme employing five controlled gates per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components.
+Although significant performance gains are typically achievable with moderate depth ($L=2$ or $3$), this inherent scaling increases the architectural footprint and component count compared to a standalone SPD.
+For instance, a three-level ESPD scheme employing five controlled modules per level ($n=5$) necessitates $6^{3}=216$ detections (in the absence of techniques such as multiplexing), resulting in hundreds of integrated components.
 While not posing a fundamental physical barrier (as discussed in Section 3.2), this resource-intensive nature introduces a non-trivial engineering challenge for immediate large-scale deployment.
-However, this complexity is well within the capabilities of modern large-scale photonic integrated circuits.
+However, this complexity is compatible with modern large-scale photonic integrated circuits.
 Recent advances have successfully demonstrated the integration of hundreds to thousands of components on single chips [54, 55, 56, 57].
-Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], verifying that the component density required for a multi-level ESPD is manufacturable using standard complementary metal-oxide-semiconductor processes.
-Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss material platforms.
+Notably, very-large-scale integration of quantum photonic components has already been experimentally demonstrated, with over 2,500 integrated on a single silicon-on-insulator chip [58], confirming that the integration density required for a multi-level ESPD is fully compatible with standard complementary metal-oxide-semiconductor processes.
+Furthermore, concerns regarding cumulative loss in cascaded logic gates can be mitigated by emerging low-loss platforms.
 For instance, silicon nitride waveguides have demonstrated propagation losses as low as 0.2 dB/cm in complex programmable processors [59], and thin-film lithium niobate platforms offer ultra-low losses of approximately 0.06 dB/cm alongside high-speed modulation capabilities [60].
 In summary, the hardware complexity of the ESPD should be evaluated as a deliberate engineering trade-off against the necessity of cryogenic operation.
-The SNSPD paradigm mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
+The SNSPD mandates bulky, energy-intensive, and expensive cryogenic infrastructure (typically operating below 4 K).
 The ESPD scheme, conversely, shifts the burden from a demanding operating environment to the domain of on-chip integration.
@@
 3.3.3 Optimization Design
-Table 7: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n,k)$. Settings are the same as those in Section 2.4
+The inherent exponential resource scaling discussed in Section 3.3.2 necessitates optimization for future practical deployments.
+Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$).
+Minimizing the iteration depth ($L$) provides significant savings in component count and system complexity due to the exponent dependence.
+As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
+Furthermore, as shown in Table 7 and visualized in Fig.10, with a suitably selected $n,k$ on each level, the convergence can be accelerated, and the intermediate-level state can be optimized, which is crucial for minimizing resource consumption.
+Table 7: Performance evolution of the ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, under varied $(n_{s},k_{s})$. Settings are the same as in Section 2.4
 Para 4 Para 5 Para 6
@@
 8 (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
-(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
-(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
-Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different $k$. Para 3 is the same as in Table.2.
-The inherent exponential resource scaling discussed in Section 3.2.2 necessitates optimization for future practical deployments.
-Specifically, the design of the ESPD requires balancing performance metrics ($DE,DCR$) against implementation complexity ($L,n,k$).
-Reducing the number of levels ($L$) by even one step can result in significant savings in component count and system complexity due to the $(n+1)^{L}$ dependence.
-As demonstrated in Section 2.4, the choice of design parameters, namely $n$ (the number of auxiliary signals per level) and $k$ (the decision threshold), plays a critical role: enlarging $k$ reduces DCR exponentially, while improving DE requires a lower $k$ and a larger $n$, with the trade-off being that enlarging $n$ increases detection cost.
-Furthermore, as shown in Table 7 and visualized in Fig.5, with a suitably selected $n,k$ on each level, the convergence speed can be accelerated, and the middle-level performance can be adjusted, which is crucial for saving costs.
+(a) DE $\eta_{s}$ vs. level $s$.
+(b) DCR $d_{s}$ vs. level $s$.
+Figure 10: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with different values of $n,k$. Para 3 is the same as in Table.2.
 The optimization problem also encompasses the optimal circuit topology.
@@
 As the controlled modules’ order can be changed and the auxiliary signals behave similarly to the control signal [^5] This justifies the dedicated filter block within the paradigm: For the controlled gate be C-NOT, the auxiliary output state is $|1\rangle$ if the input signal is non-vacuum, while is $|\emptyset\rangle$ if the input signal is vacuum, theoretically consistent to the input signal after state-preparation., alternative designs could involve leveraging a binary-tree structure and defining the positive report event via a global decision criterion, rather than the current level-by-level $n$-length structure with local (level-wise) criterion used in Section 2.1.
-However, these architectural changes introduce new theoretical challenges: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
+However, these architectural changes introduce new theoretical complexity: the general binary-tree structure may result in non-independent distributions that are difficult to analyze, and the global decision criterion selection must carefully account for both device requirements and implementation convenience.
 3.3.4 Physical Integrations and Experiments
 Despite the convincing evidence provided by the numerical simulations in Section 2.4, the physical implementation of the ESPD scheme still requires dedicated experimental validation.
-Reassuringly, as established in Section 3.1, the scheme relies only on readily available conventional quantum optical components, suggesting no fundamental technological barriers are preventing its realization.
-However, the realization of the large-scale iterative circuit demands significant and focused engineering efforts, particularly concerning the precise integration and synchronization of the cascaded components.
-Furthermore, the seamless integration of the enhanced detector into various quantum optical tasks (such as QKD) needs full investigation.
+Reassuringly, as established in Section 3.2, the scheme relies only on readily available quantum optical components, suggesting no fundamental technological barriers prevent its realization.
+However, the realization of the large-scale iterative circuit demands significant engineering efforts, particularly concerning the precise integration and synchronization of components.
+Furthermore, the seamless integration of the ESPD into various quantum optical tasks (such as QKD) needs full investigation.
 Thus, both the physical realization of the ESPD scheme and the characterization of its performance in applied scenarios are key directions for future work.
@@
 While the ESPD framework significantly enhances DE and suppresses DCR, it generally inherits the timing characteristics of the baseline detectors.
-Specifically, the system’s timing jitter and maximum count rate remain constrained by the intrinsic performance of the initial SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices).
-However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, ensuring negligible added temporal broadening.
+Specifically, the system’s timing jitter and count rate are primarily determined by the intrinsic temporal response of the baseline SPDs (e.g., the avalanche buildup time and dead time of InGaAs/InP devices).
+However, the ESPD architecture does not introduce significant additional penalties in principle: the optical-controlled quantum operations function on timescales usually faster than the electronic response of room-temperature SPDs, such that the architectural integration introduces negligible temporal broadening.
 Furthermore, as the final SPDs operate in parallel, the effective system throughput remains comparable to that of the constituent devices.
-Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics.
+Nevertheless, future iterations of the ESPD scheme could incorporate techniques such as multiplexing strategies to optimize these temporal metrics as well.
 4 Conclusion
-In summary, the enhanced single-photon detection (ESPD) framework is established to reformulate photon detection as an iteratively enhanced quantum-information-processing task composed of state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a fixed device-level operation.
-Within this system-level theoretical framework, detection performance is progressively improved through structured quantum processing using exclusively non-superconducting, room-temperature components.
-Numerical simulations, grounded in experimentally reported parameters and commercially available photonic components, demonstrate that the ESPD framework can substantially enhance the effective detection efficiency (DE) (from 59% to 93%) while strongly suppressing the effective dark count rate (DCR) (from $10^{-2}$ to $10^{-9}$) even when initialized with a conventional room-temperature SPD reported two decades ago.
-The resulting effective performance metrics are comparable to those of state-of-the-art superconducting single-photon detectors, significantly relaxing the minimal tolerable channel transmission rate in quantum communication applications.
-Beyond performance enhancement, the ESPD framework makes explicit the fundamental trade-off between resource overhead and operating conditions, replacing cryogenic infrastructure requirements with system-level integration complexity.
-Although the realization of large-scale ESPD architectures would require further advances in device integration and experimental implementation, the theoretical formulation and parameter studies presented here establish a concrete foundation for the framework and provide concrete guidance for future experimental efforts.
-In essence, this work establishes a conceptually complete and theoretically grounded framework for high-performance single-photon detection at room temperature.
-It demonstrates that intrinsic device-level limitations need not be overcome solely through component replacement, but can instead be effectively compensated by system-level architectural design.
-As such, the ESPD framework offers a viable pathway toward next-generation photon detection and provides a general methodology for rethinking performance limits in quantum technologies under practical constraints.
-Acknowledgements
-The author acknowledges the use of Gemini and ChatGPT (accessed December 2025) to, and only to, support language polishing, flow paragraphs, and correct grammar, to improve readability.
-All AI-generated suggestions were reviewed, revised, and approved by the authors, who take full responsibility for the accuracy and integrity of the work.
-All insight, results, methods, analysis, and experiments are ordinarily created and owned by the author.
+In summary, we have established the enhanced single-photon detection (ESPD) framework, which reconceptualizes photon detection as an iteratively enhanced quantum-information-processing task that integrates state preparation, controlled operations, projective measurements, and multi-copy decision analysis, rather than as a static device-level operation.
+Within this system-level theoretical framework, the detection performance is formulated as a non-linear dynamical system, where the effective metrics are progressively optimized through structured quantum processing using exclusively non-superconducting, room-temperature components.
+Analytical approximations, Monte Carlo analysis, and numerical simulations, grounded in experimentally reported parameters and commercial photonic specifications, demonstrate that the ESPD framework drives detection performance toward the high-performance basin of attraction, provided the initial conditions lie within a suitable operating range.
+Consequently, the performance of traditional single-photon detectors (SPDs) can be substantially enhanced: the effective detection efficiency (DE) can be boosted (e.g., from 59% to 93%) while the effective dark count rate (DCR) can be suppressed by several orders of magnitude (e.g., from $10^{-2}$ to $10^{-9}$), to the level comparable to state-of-the-art superconducting SPDs, resulting in the significantly expansion of the operational boundaries of quantum communication by relaxing the minimal tolerable channel transmission rate.
+Beyond performance enhancement, the ESPD framework establishes a trade-off between resource overhead and environmental constraints, effectively translating the demand for cryogenic cooling during operation into a one-time integration complexity.
+While large-scale implementation poses engineering challenges in component synchronization and integration density, the presenting theoretical formulations and technological feasibility studies establish a rigorous foundation and provide a roadmap for future experimental realizations.
+In essence, this work provides a theoretically grounded framework for high-performance photon detection at room temperature.
+It demonstrates that intrinsic device-level limitations can be circumvented through system-level architectural design rather than reliance exclusively on material breakthroughs.
+By establishing the dynamical evolution of the ESPD paradigm, this work not only offers a viable pathway toward next-generation room-temperature photonic devices but also introduces a general methodology for rethinking performance limits in quantum technologies under real-world constraints, potentially influencing a wide range of quantum applications.
 Code Availability